### v1

# Channel-Level Calibration Methods of Silicon Photomultiplier for JUNO-TAO Central Detector

Abstract
The Taishan Antineutrino Observatory (TAO or JUNO-TAO) is a satellite observatory for the Jiangmen Underground Neutrino Observatory (JUNO), located 44 meters away from the No.1 reactor of the Taishan Nuclear Power Plant. TAO can measure the reactor antineutrino energy spectrum with excellent energy resolution (better than 2% at 1 MeV) using state-of-the-art Silicon Photomultipliers (SiPMs) operated at low temperature. To achieve this goal, the SiPMs (together with their readout electronics) must be well calibrated. This paper presents the channel-level calibration methods for the dark count rate (DCR), relative photon detection efficiency (PDE), time offset, gain, and internal optical crosstalk (IOCT) of the SiPMs based on charge and time information of the collected events. For the calibration of the external optical crosstalk (EOCT), in terms of its rate and emission angle distribution, a novel method is proposed by switching on and off different groups of SiPMs with an LED placed in the detector. Using one million simulated events, the expected calibration biases are evaluated for all the aforementioned parameters: relative PDE (2.1%), IOCT (1.4%), DCR (0.4%), EOCT Rate ($0.1\%$), gain ($0.1\%$), time offset (0.027 ns). The emission angle distribution of the EOCT photons could be measured with a bias of less than 4% in main angular range.
1 Introduction
Before the precise measurements of the reactor neutrino energy spectrum by experiments such as Daya Bay [Ref1], RENO [Ref2], and Double Chooz [Ref3], the primary theoretical models for calculating the expected reactor neutrino energy spectrum were the Summation model [Ref4] based on nuclear database data [Ref33, Ref34, Ref35, Ref36, Ref37, Ref38] and the $\beta$ conversion method such as Huber-Muller model [Ref5, Ref6] based on the ILL High Flux Reactor measurement data [Ref7, Ref8, Ref9]. However, according to the measurement results from the Daya Bay, RENO, and Double Chooz experiments, both the Summation model and the Huber-Muller model exhibit systematic deviations from the experimental results in terms of the predicted energy spectrum flux and the shape of the energy spectrum at 5 MeV. Although Estienne and Fallot have improved the Summation model [Ref10]; and Kopeikin, Skorokhvatov, Titov [Ref11]; Hayen, Kostensalo, Severijns, Suhonen [Ref12]; Giunti, Li, Ternes, Xin [Ref13] have refined the Huber-Muller model, these models currently exists that can not simultaneously explain the systematic deviations in both flux and spectrum shape. Moreover, none of models can reliably predict the fine structures in the reactor antineutrino spectrum [Ref46]. If the JUNO experiment directly uses these predicted spectra as inputs, it will introduce significant uncertainties in the energy spectrum, which will affect the sensitivity of the neutrino mass ordering measurement [Ref14]. To reduce the model dependence of the reactor neutrino energy spectrum, JUNO has deployed a satellite detector, the Taishan Antineutrino Observatory (TAO or JUNO-TAO) at a baseline of 44 meters away from Unit 1 of the Taishan Nuclear Power Plant [Ref15].
[…]
To achieve the aforementioned physical goals, it is crucial to attain the designed energy resolution for the TAO experiment. However, the negative effects of SiPMs can directly impact the energy resolution. For instance, according to the TAO Conceptual Design Report, DN and optical crosstalk (OCT) contribute to a degradation of 0.75% and 0.5% on the energy resolution at 1 MeV [Ref15]. In addition to affecting the energy resolution, these two effects can cause deviations of 1 mm and 2 mm on the vertex reconstruction, as reported by Shi [Ref16]. Therefore, accurate calibration of all relevant SiPM parameters for each readout channel is of crucial importance.
[…]
This paper investigates and presents channel-level calibration methods for SiPMs installed in the TAO CD, including calibration of the dark count rate (DCR), relative photon detection efficiency (PDE), internal optical crosstalk (IOCT) rate, gain, and time offset. Additionally, a method based on selectively turning on and off different groups of SiPMs utilizing the fine control capability of the SiPM high-voltage system and an LED calibration source is proposed to calibrate the EOCT rate and emission angle distribution. This study is conducted using simulated data generated by the TAO offline software.
This paper is structured as follows: Section 2 introduces the TAO detector; Section 3 presents the simulated model; Section 4 describes the calibration of the DCR, relative PDE, and time offset based on hit time information; Section 5 details the calibration of gain and IOCT rate based on charge information; Section 6 introduces a novel method for EOCT rate and emission angle distribution calibration; Section 7 summarizes the results and discusses limitations of these methods.
2 TAO Detector
The TAO detector structure is shown in Figure 1. It consists of three sub-detectors: the central detector, water tanks and the top veto tracker.
Figure 1: Design of the TAO detector
The central detector contains 2.8 tonnes of GdLS [Ref20] enclosed within an acrylic sphere. The exterior of the acrylic sphere is covered by a copper shell, which inner surface is equipped with 4024 HPK SiPMs, each measuring $50.7\times 50.7$ ${\mathrm{mm}}^{2}$. The copper shell is housed inside a stainless steel tank filled with linear alkylbenzene, which serves as a buffer liquid to shield external radioactive background, maintain temperature stability, and provide optical coupling between the acrylic container and the SiPMs. Both the tank and the copper shell have refrigeration tubes that cool the central detector to -50 °C to reduce the SiPM DCR to below 100 ${\mathrm{mm}}^{2}$. At the top of the CD is the calibration system [Ref21], which includes an Automatic Calibration Unit (ACU) and a Cable Loop System (CLS). The ACU can deploy calibration sources (include LED source, $^{68}{\rm Ge}$ source and combined $\gamma$ sources) along the central axis of the GdLS. It is used to calibrate the CD’s non-linearity and perform channel-level calibration of SiPMs. The CLS consists of the stainless steel cable passing through specific anchor points with a small segment plated with $^{137}{\rm Cs}$, which is used to calibrate the non-uniformity of the central detector.
Surrounding the central detector are three water tanks. Each tank is filled with 50 tonnes of pure water, which can shields external radioactive background and produces water Cherenkov light signals when muons pass through. The tanks are equipped with 300 3-inch PMTs, which detect the water Cherenkov signals to veto muons.
Above the central detector is the top veto tracker [Ref47, Ref48], which consists of four layers of plastic scintillators and is designed for muon veto.
3 TAO Central Detector Simulated Model
This study utilizes simulated data generated by the TAO offline software to investigate calibration methods for SiPM parameters. This section presents the simulation model.
TAO offline software is a tool for experimental data processing, physics data analysis and simulated data production. As a key component of the TAO experiment, the TAO offline software is built upon the SNiPER [Ref23] framework and includes modules for physics event generator, detector simulation, electronics simulation, raw data conversion, calibration, reconstruction, and physical analysis. For simulated data production, the physics event generator, detector simulation and electronics simulation modules are utilized.
3.1 Physics Event Generator
The physics event generator is used to produce particles and specific events, and the physics event generator of TAO directly adopts that of JUNO. In JUNO simulation paper, it indicates that we have three types physics event generators: particle gun, HepEvt and GENIE [Ref24]. In this paper, we use particle gun which is a tool used to generate particles with specific particle types, momenta and positions.
3.2 Detector Simulation
Following the physics event generator, the detector simulation is performed. The TAO detector simulation is based on Geant4 (version:10.04.p02) [Ref41], and models the propagation of particles through the detector. The TAO detector simulation module is largely derived from the Daya Bay detector simulation and has been modified in the following four aspects:
  1. 1.
    Detector geometry construction: The detector geometry has been changed to TAO’s [Ref15].
  2. 2.
    Physics processes: The physics process has been changed to JUNO’s [Ref24].
  3. 3.
    Optical model of the LS: The optical model is constructed based on three components: linear alkylbenzene (LAB), 2,5-diphenyloxazole (PPO) and p-bis-(o-methylstyryl)-benzene (bis-MSB) [Ref49]. The absorption-remission spectrum is determined by PPO and bis-MSB [Ref50], and the light attenuation length is jointly determined by the absorption probabilities of LAB, PPO, and bis-MSB. Additionally, the Birks constant has been updated to match that used in the JUNO simulation.
  4. 4.
    Sensitive detector: The sensitive detector has been changed from PMTs to SiPMs. The photon detection efficiency and reflectivity have been configured according to experimental measurements, where the photon detection efficiency is about 50% and the reflectivity is about 8.5%. [Ref32, Ref51].
3.3 Electronics Simulation
The electronics simulation module is designed to simulate the process from SiPM photoelectrons to the readout of electrical signals by the trigger system. It primarily consists of three components: SiPM effect modeling, event triggering, and waveform simulation.
3.3.1 SiPM Effect Modeling
When events from the detector simulation are input into the electronics simulation module, each hit in each event is converted into an electronic pulse. The electronics simulation module add SiPM effects, which including dark noise (DN), gain, time offset, optical crosstalk (OCT), and afterpulse (AP).
[…]
The gain determines the amplitude of each pulse. The amplitude of each pulse is sampled from a Gaussian distribution with a mean value of 1 and the standard deviation of 15% corresponding to the charge resolution of electronics system [Ref15]. The amplitude of the pulse indicates how many times greater the waveform amplitude is that of a single photoelectron (SPE) waveform. By integrating the SPE waveform, the true gain value can be determined which is 3000 ADC counts.
Regarding the time offset, the timestamp of each pulse is adjusted by adding a constant time offset specific to the corresponding channel. In this study, a 20 ns time offset is applied to the SiPMs located in the northern hemisphere. Additionally, a Gaussian smearing with a standard deviation of 1 ns [Ref15] is added to each pulse timestamp for time spread.
For the OCT, each pulse can generate OCT pulses (including the OCT itself), and the number of OCT pulses generated by each pulse is determined via Poisson sampling. OCT is divided into internal optical crosstalk (IOCT) and external optical crosstalk (EOCT). The IOCT refers to OCT pulses occurring in the same channel as the source pulse, while the EOCT occurs in different channels. In the TAO electronics simulation, IOCT and EOCT sampling are independent, each controlled by separate Poisson parameters, and every pulse undergoes both IOCT and EOCT sampling. During the TAO SiPM mass test, the SiPMs were tested in air; however, the OCT rates in air and LAB are different. Therefore, the OCT parameters used in the simulation are not from the TAO SiPM mass test, they are from Guan’s paper, which measured the SiPM OCT rate in the LAB [Ref17]. However, in his paper, the IOCT rate is very low, making it unsuitable for studying the IOCT calibration method. Consequently, we set the IOCT Poisson parameter to 0.15. The EOCT Poisson value is 0.2447 multiplied by the detector geometric factor, which corrects for the fact that the SiPM coverage is not 100%. The detector geometric factors vary depending on each channel, with values ranging from 0.7 to 0.8.
For the AP, each main pulse may generate an AP. The occurrence of an AP is determined by sampling from a Bernoulli distribution with an AP rate of 1%, as obtained from the TAO SiPM mass test [Ref32]. The delay of the AP relative to the main pulse follows an exponential distribution with a parameter of 50 ns. The delay time interval between the AP and the main pulse determines the amplitude of the AP, which is given by the expression $(1-e^{\frac{-\Delta t}{\tau}})$, where $\Delta t$ is the time interval and $\tau$ is the recovery time of the SiPM, set to 32.5 ns. These two time constants are from Rolph’s paper [Ref52].
The parameters are summarized in Table 1. The typical ranges for DCR, PDE and AP are based on the TAO SiPM mass test result [Ref32]. The values for charge resolution and time spread are taken from our conceptual design report [Ref15]. The typical ranges for internal and external OCT are sourced from Guan’s paper [Ref17]. The gain value is derived from our prototype detector result [Ref53].
Table 1: Summary of the simulated data used for the channel-level calibration of the SiPM parameters
Parameter Typical Range Simulated
[…]
Gain 3000 ADC Count to 5000 ADC Count 3000 ADC Cout
Charge Resolution  15% 15%
Time Offset Unknown 0 ns or 20 ns
Time Spread  1 ns 1 ns
PDE 45% to 55% 50%
Internal Crosstalk < 1% 15%
External Crosstalk 20% to 30% 24.47%
Afterpulse < 1% 1%
3.3.2 Event Triggering
After converting the SiPM hits into electronic pulses and adding the pulses from SiPM effect, we can perform event triggering. The TAO electronics support two trigger modes: the nhit self-trigger and the external trigger. The nhit self-trigger scans a 300 ns time window along the time axis and triggers when the number of pulses within this window exceeds a threshold of 1000; the left boundary of the window is designated as the trigger time. The external trigger is forced to trigger at the start time of the detector simulation event.
3.3.3 Waveform Simulation
After triggering, the pulses from the trigger event are converted into waveforms. Each pulse is converted into a SPE waveform with the waveform starting point determined by the the pulse’s timestamp, and the waveform amplitude corresponding to the pulse’s amplitude. Subsequently, white noise is added with an amplitude equal to 10% of the SPE waveform amplitude.
After generating the waveforms, the charge and time of each waveform can be calculated. When three consecutive points exceed the threshold (set at 50% of the SPE amplitude), the waveform time is determined by the intersection of the line (fitted to the rising edge of the waveform) and the baseline. The charge is obtained by integrating the waveform backward from the threshold-crossing point to the baseline and forward to the baseline.
3.4 Simulated Data Sets
The simulated data sets used for SiPM calibration in this study are listed in Table 2. The DCR, IOCT and gain can be calibrated using arbitrary event types. In this study, we use electrons uniformly distributed within the LS with energies following a uniform distribution in the range of 1–10 MeV to represent arbitrary events. For the time offset calibration, the calibration source must be located at the center of the LS. This placement ensures that photons arrive simultaneously at all the SiPMs, thereby eliminating systematic bias induced by photon flight time. Calibration of the relative PDE requires that the incident photon flux on each SiPM is statistically uniform. Therefore, the calibration source is deployed at the center of the LS to ensure that the light field is isotropic. The $^{68}{\rm Ge}$ source fulfills this criterion, as it produces oppositely directed $\gamma$-ray pairs from positron annihilation, and the light emission from the LS is isotropic. For the calibration of EOCT, the SiPM-switching method is adopted, which requires the use of LED-based external triggering. The specifics of this methodology are elaborated in Section 6.
In the simulated data, the true values of all parameters are known. Accordingly, calibration accuracy is quantified as the bias and standard deviation. The calibrated bias is calculated as the average difference between the calibrated and true simulated values across all channels, and the calibrated standard deviation is obtained through the standard deviation of each channel calibrated value after subtracting the corresponding simulated true value (excluding the bias of relative PDE).
Table 2: Summary of the simulated data used for the channel-level calibration of the SiPM parameters
Parameter Particle Position Trigger Type Data Type True Value
[…]
Gain Arbitrary Arbitrary Self Trigger Charge 3000 ADC Count
Time Offset LED LS Centre External Trigger Hit Time 0ns or 20ns
Relative PDE $^{68}{\rm Ge}$ LS Centre Self Trigger Hit Time 1
Internal Crosstalk Arbitrary Arbitrary Self Trigger Charge 0.15
External Crosstalk LED LS Centre External Trigger Charge From EOCT disabled data
4 Calibration Based on Hit Time Information
4.1 Dark Count Rate
[…]
The DCR quantifies the level of DN of SiPM and is defined as the number of DN generated per unit sensitive area of the SiPM per unit time. It can be estimated by counting the hits occurring in the plateau before the peak of the main hits generated by the physics events, as shown in Figure 2, which displays the simulated hit time distribution of a channel. With $n$ denoted as the total number of the counted hits in the plateau (indicated by the red line in the figure), $N$ as the number of physics events, $t$ as the duration of the plateau before the trigger timestamp (from -100 ns to 0 ns), and $S$ as the sensitive area of the SiPM. Then, the DCR can be calculated using Equation 4.1:
$DCR=\frac{n}{N\cdot t\cdot S}$ (4.1)
Figure 2: The Simulated Hit Time Distribution for a Channel. The trigger algorithm slides a 300 ns time window along the time axis. An event is triggered when the number of channels with hit within this trigger time window exceeds a preset threshold (set to 1000 in the TAO electronics simulation). The trigger time corresponds to the start of the trigger time window and is aligned to t = 0 ns. The scintillation light from the physics event is emitted within a short time after the trigger time, so hits occurring before the trigger time primarily originate from DN.
However, the hits observed in the plateau range may also include minor contributions from hits generated by the EOCT of DN from other SiPMs, as well as hits from delayed APs. For the TAO SiPM, the delayed AP has been measured to be very weak [Ref26]. Therefore, only EOCT contamination needs to be considered in the DCR calibration.
The DCR calibration result is shown in the left part of Figure 3. Since the SiPMs are mounted in circle around the central $z$-axis of the detector and layer by layer along the central $z$-axis, the channels have azimuth number and layer number as its coordinate. A significant bias of 23.6% with the standard deviation of 1.7%. However, when the EOCT effect is disabled, the bias reduces is substantially reduced to -0.2% as presented in the right part of Figure 3. This clearly demonstrates that the EOCT effect introduces a considerable bias in the calibration result. A correction method will be presented in Section 6.3.
[…]
4.2 Time Offset
The hit time of the SiPM obtained via the electronic readout satisfies Equation 4.2, where $t_{R}$ is the hit time from the electronic readout. It is calculated as $t_{R}=t_{A}-t_{T}$, where $t_{A}$ is the absolute timestamp and $t_{T}$ is the trigger timestamp. The absolute timestamp consists of several components: $t_{P}$ is the time when the particle reaches and deposits energy in the LS; $t_{\rm LS}$ is the time when the LS emits scintillation photons; $t_{\rm tof}$ is the time of flight of scintillation photons to the SiPM; the electronic processing time, which is divided into two parts: $t_{\rm same}$ and $t_{\rm timeoffset}$. Here, $t_{\rm same}$ represents the identical portion of the processing time across all electronic channels, while the differing portions constitute the time offset ($t_{\rm timeoffset}$) in different channels.
$\begin{split}t_{R}&=t_{A}-t_{T}\\ t_{A}&=t_{P}+t_{\rm LS}+t_{\rm tof}+t_{\rm same}+t_{\rm timeoffset}\\ t_{\rm timeoffset}&=t_{R}+t_{T}-t_{P}-t_{\rm LS}-t_{\rm tof}-t_{\rm same}\end{split}$ (4.2)
Since the TAO experiment only requires correcting the relative time offsets between electronic channels [Ref15] to ensure that the timestamps of all channels are synchronized, this study deploys an LED source at the center of the LS to calibrate the relative time offset between different channels. The LED source acquires events using an external trigger method, ensuring that the trigger time $t_{T}$ is identical across all channels. Therefore, $t_{T}$ does not affect the calibration of relative time offset. The $t_{P}$ and $t_{\rm LS}$ are statistically identical across all channels, so the $t_{P}$ and $t_{\rm LS}$ does not influence relative time offset calibration. Because the LED is deployed at the center of the LS, the distance from the LS center to each SiPM channel is the same, meaning the $t_{\rm tof}$ is identical across all channels. Consequently, the relative time offsets between different channels can be derived from the relative $t_{R}$ differences for each channel.
As shown in Figure 5, we use the double-sided Crystal Ball function to fit the distribution of $t_{R}$. The mean value of the fitted function corresponds to the $t_{R}$ value for the channel. When a reference channel is selected, the time offset of $\rm channel_{\rm i}$ relative to the reference channel ($t^{\rm chid=i}_{\rm timeoffset}$) is the readout time of $\rm channel_{\rm i}$ ($t^{\rm chid=i}_{R}$) minus the readout time of the reference channel ($t^{\rm RefCh}_{R}$).
$t^{\rm chid=i}_{\rm timeoffset}=t^{\rm chid=i}_{R}-t^{\rm RefCh}_{R}$ (4.3)
Figure 4: Double-Sided Crystal Ball Function Fit to the Time Spectrum.
The time offset calibration result is shown in Figure 5, the bias is 0.027 ns and the standard deviation is 0.031 ns.
Figure 5: Time Offset Calibration Result. The left figure shows the calibration result for each channel. In this simulation, half of the channels are set to the time offset of 0 ns, while the other half are set to the time offset of 20 ns. The middle and right figures show the bias between the time offset calibration result and the mock truth.
In the actual TAO CD, the the calibration source’s position may be offset from its deployed position, and this offset contributes to a bias in the time of flight. Specifically, a 1 cm offset results in a 0.13 ns bias between the SiPMs located at the south and north poles of the CD. Considering that the time resolution of the TAO electronics exceeds 1 ns [Ref15], both the calibration bias and the calibration source’s positional offset from the LS center are acceptable.
4.3 Relative Photon Detection Efficiency
As shown in Equation 4.4, the photon detection efficiency (PDE) of a SiPM is defined as the ratio of the number of photons detected by the SiPM ($N_{\rm Det}$) to the number of photons incident on the SiPM surface ($N_{\text{photon}}$). The PDE is determined by the SiPM’s quantum efficiency (QE), electron avalanche probability, and geometric fill factor [Ref27, Ref28]. Specifically, the QE depends on the wavelength of the photon, the electron avalanche probability is a function of the operating voltage of the SiPM, and the geometric fill factor is constant for each SiPM.
$PDE=\frac{N_{\rm Det}}{N_{\text{photon}}}$ (4.4)
Since the number of photons incident on the SiPM surface depends on the particle type, particle energy, light yield of the LS, and light attenuation length of the LS, calibrating the absolute PDE of each SiPM is impractical. However, calibrating the relative PDE is feasible. The TAO calibration system employs three types of calibration sources: an LED source, a $^{68}{\rm Ge}$ source and a combined $\gamma$ source [Ref21]. The light field produced by LED source may be not isotropic, and the combined $\gamma$ source consists of multiple $\gamma$ sources with different energies. In contrast, the $^{68}{\rm Ge}$ source is a positron annihilation $\gamma$ source. Due to the presence of the source shell, positrons annihilate inside the shell, producing a pair of $\gamma$ emitted in opposite directions, which enhances the uniformity of the light field. Therefore, the $^{68}{\rm Ge}$ source is used for relative PDE calibration.
Deploying the $^{68}{\rm Ge}$ source at the center of the LS, the number of photons detected by each channel follows a Poisson distribution. The parameter of the Poisson distribution is $N_{\rm Det}$ which represents the mean number of detected photons. It satisfies Equation 4.2 where $N_{\text{PE}=0}$ is the number of events with zero hits in the channel, and $N_{\text{Event}}$ is the number of events. Since the $^{68}{\rm Ge}$ source is deployed at the center of the LS and the light emission from the LS is isotropic, the average number of incident photons on the surface of each channel is the same. According to Equation 4.3, the relative PDE of each channel corresponds to the relative $N_{Det}$ value.
$N_{\rm Det}=-ln(\frac{N_{\text{PE}=0}}{N_{\text{Event}}})$ (4.5)
However, the DN effect can also generate hits. If the DCR of different channels vary, the values of $N_{\text{PE}=0}$ will differ, causing biases in the relative PDE calculated based on the relative $N_{\rm Det}$. Therefore, it is necessary to correct $N_{\text{PE}=0}$ before calculating the relative $N_{\rm Det}$. As shown in Figure 2, after correcting for the time offset effect, we count the hits in the time range from 0 ns to 160 ns, considering them as contributions from DN, and denote the number of events with hits in this DN time range as $N_{\rm DN}$. We also count the hits in the time range from 240 ns to 400 ns, considering them as contributions from both DN and $^{68}{\rm Ge}$, and denote the number of events with hits in the $^{68}{\rm Ge}$ time range as $N_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$. As shown in Equation 4.6, $P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$ is the probability of hits generated by DN, and $P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$ is the probability of hits generated by both DN and $^{68}{\rm Ge}$. Since DN and $^{68}{\rm Ge}$ are independent, the combined probability is the sum of the probabilities of hits generated by DN alone, by $^{68}{\rm Ge}$ alone, and by both simultaneously. Therefore, the probability of hits generated by Ge68 $P_{\rm{}^{68}{\rm Ge}}$ can be derived from $P_{\rm DN}$ and $P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$. In this simulation, we set four different DCR values: 20, 40, 60, and 80 ${\mathrm{mm}}^{2}$. The results before and after applying this correction method are shown in Figure 6, this method can effectively correct the bias caused by different DCR levels.
$\begin{split}P_{\rm DN}&=\frac{N_{\rm DN}}{N_{\text{Event}}}\\ P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}&=\frac{N_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}}{N_{\text{Event}}}\\ &=P_{\rm DN}\cdot(1-P_{\rm{}^{68}{\rm Ge}})+P_{\rm{}^{68}{\rm Ge}}\cdot(1-P_{\rm DN})+P_{\rm{}^{68}{\rm Ge}}\cdot P_{\rm DN}\\ P_{\rm{}^{68}{\rm Ge}}&=\frac{P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}-P_{\rm DN}}{1-P_{\rm DN}}\\ N_{\text{PE}=0}&=(1-P_{\rm{}^{68}{\rm Ge}})\cdot N_{\text{Event}}\end{split}$ (4.6)
Figure 6: Calibration Bias Before and After DCR Correction. The left figure shows the bias which uncorrected DN differences, and the right figure shows bias which corrected.
After Correcting the DN differences in each channel, the relative PDE calibration result is shown in the left part of Figure 7. There is an approximate 3% bias between the channels at the pole and equator, which is caused by optical reflection on the SiPM surface. As shown in the right part of Figure 7, this bias disappears after disabling the optical reflection on the SiPM surface in the detector simulation. This bias will be corrected based on the result of the detector optical model and simulation in the future.
Figure 7: Bias Between Relative PDE Calibration Result and Mock Truth. The left figure shows the results under normal configuration, while the right figure shows those with the optical reflection effect on SiPM surface disabled.
5 Calibration Based on Charge Information
5.1 Gain
The gain is quantified as the average charge output per single photoelectron (PE) avalanche in the SiPM after amplification by the front-end electronic board [Ref29]. Consequently, the gain depends on the operating voltages of both the SiPM and the front-end board. In the TAO CD, these two voltages are stable.
$\begin{split}f=\sum_{i=1}^{n}N_{i}\cdot Gaus(\mu_{i},\sigma_{i})\end{split}$ (5.1)
Figure 8: Multi-Gaussian Function Fit to Charge Distribution. The horizontal axis is measured in ADC counts, where 1.8 volts correspond to $2^{16}$ ADC counts. Each peak represents the charge response for a specific number of PE: the first peak corresponds to 1PE, the second to 2 PEs, and so on. The separation between adjacent PE peaks indicates the gain.
As shown in Figure 8, the charge generated by a single PE follows a Gaussian distribution, and the individual photons detected by the avalanche photodiodes (APDs) in the SiPM are independent. Consequently, the total charge resulting from the simultaneous detection of multiple PEs also follows a Gaussian distribution, and the overall distribution is a multi-Gaussian distribution (Equation 5.1 where f is the multi-Gaussian function, $N_{i}$ is the number of hits corresponding to each PE, $\mu_{i}$ is the mean value, and $\sigma_{i}$ is the standard deviation). Different Gaussian peaks correspond to the charge responses for different numbers of PEs. By fitting the charge spectrum with a multi-Gaussian function, the gain is determined from the separation between adjacent peaks.
The gain calibration result is shown in Figure 9. The unit of gain is ADC counts, where 1.8 volts corresponds to $2^{16}$ ADC counts. The bias is 0.084% and the standard deviation is 0.094%.
Figure 9: Bias Between Gain Calibration Result and Mock Truth.
5.2 Internal Optical Crosstalk
When an electron undergoes an avalanche process in the avalanche region of the APD, the avalanche electron accelerates in the electric field and can emit photons. These photons have a probability of escaping from the APD and being detected by other APDs. This phenomenon is known as the optical crosstalk (OCT) effect in SiPMs [Ref44, Ref30]. Specifically, if an APD generates an OCT photon and this photon is detected by another APD within the same electronic channel, it is defined as internal optical crosstalk (IOCT). In contrast, if the APD detecting the OCT photon is in a different channel, it is referred to as external optical crosstalk (EOCT). Therefore, the OCT effect depends on the electric field intensity inside the APD, which is a function of the SiPM operating voltage. In TAO CD the operating voltage of the SiPM is stable.
There are two methods for calibrating the IOCT in the TAO CD: multiple photoelectrons (multi-PEs) hit analysis and generalized Poisson fitting. For multi-PEs hit analysis, the calibrated quantity $P_{\rm IOCT}$ represents the probability that a single electrical pulse generates IOCT pulses. In generalized Poisson fitting, the calibrated quantity $\lambda$ is the branching Poisson distribution parameter of the IOCT. It means that one pulse generates an average of $\lambda$ IOCT pulses following a Poisson distribution, which constitutes the first Poisson branch. Subsequently, each IOCT pulse in the first branch also generates average of $\lambda$ IOCT pulses, forming the second branch and subsequent branches follow similarly. The total distribution of these Poisson branches is branching Poisson distribution [Ref31]. The conversion relationship between these two quantities is given by Equation 5.2.
$P_{\rm IOCT}=1-exp(-\lambda)$ (5.2)
5.2.1 Multiple Photoelectron Hit Analysis
Hits occurring before the trigger time are generated by DN pulses and EOCT pulses induced by DN, which are referred to as equivalent dark noise (EDN) in this paper. These EDN pulses can induce IOCT. When IOCT is induced by these pulses, a multi-PEs hit composed of both the EDN and IOCT signals is invariably generated. In contrast, the probability of forming a multi-PE hit from EDN alone is minimal. As shown in Figure 10, it can be observed that multi-PEs hits are primarily contributed by the IOCT effect.
Figure 10: Charge Spectrum Distribution with and without IOCT. The red line represents a case with IOCT ($\lambda$ = 0.15), and the blue line represents a case without IOCT. It can be observed that the multi-PEs component is mainly contributed by IOCT.
Therefore, the multi-PEs region of the charge spectrum acquired in a dark environment is dominated by IOCT. Hence, the IOCT rate can be quantified using Equation 5.3, where $P_{\rm IOCT}$ is the probability of IOCT occurrence, $N_{\rm multPE}$ is the number of multi-PEs hits and $N_{\rm total}$ is the total number of hits.
$P_{\rm IOCT}=\frac{N_{\rm multPE}}{N_{\rm total}}$ (5.3)
Figure 11: Bias Between IOCT Calibration Result and Mock Truth. The parameter calibrated for IOCT corresponds to the probability of IOCT occurrence.
The IOCT calibration result is shown in Figure 11, the bias is 5.70% and the standard deviation is 2.73%. By analyzing the simulated data, which close IOCT and AP effect, the multi-PEs count rate is 0.74%. Therefore, a portion of the bias is attributable to the multi-PEs hits induced by EDN. The bias contribution is quantified by Equation 5.4, where $P_{\rm IOCT}$ is the probability of IOCT occurrence, $P_{\rm EDNM}$ is the probability that multi-PE hits are generated by EDN, $P_{\rm AP}$ is the probability of AP occurrence. The bias contributed by multi-PEs EDN is 4.57%. When the IOCT effect is disabled and 1% AP effect is added in the simulated data, the multi-PEs hit rate from 1 PE EDN hits is 0.23%, corresponding to a bias is 1.42%. It is important to note that since the IOCT calibration relies exclusively on DN hits, the limited statistics of the DN sample result in a relatively large standard deviation in the calibration results.
$\begin{split}Bias_{\rm EDNM}&=\frac{(1-P_{\rm IOCT})\cdot P_{\rm EDNM}}{P_{\rm IOCT}}\\ Bias_{\rm APM}&=\frac{(1-P_{\rm IOCT})\cdot P_{\rm AP}}{P_{\rm IOCT}}\end{split}$ (5.4)
5.2.2 Generalized Poisson Fitting
To eliminate the influence of multi-PEs EDN on the IOCT calibration, the charge spectrum can be fitted using a generalized Poisson distribution. As demonstrated in Vinogradov’s paper [Ref31], when $\mu$ primary pulses cause IOCT, the number of the first branch IOCT pulses generated by each primary pulse follows a Poisson distribution with a parameter $\lambda$. Each pulse in the first branch of IOCT can further generate the second branch of IOCT also following the same Poisson distribution; this pattern continues for subsequent branches. The total distribution of the number of pulses from the $\mu$ primary pulses and IOCT follows a the generalized Poisson distribution. Since IOCT pulses have no delay relative to the primary pulses, the distribution of the number of pulses corresponds to the number of PEs. The function of the generalized Poisson distribution function is shown in Equation 5.5, where k is the number of PE, $\mu$ is the number of primary pulse, and $\lambda$ is the branching Poisson distribution parameter of the IOCT.
$GP(k,\mu,\lambda)=\frac{\mu\cdot(\mu+\lambda\cdot k)^{k-1}\cdot exp(-\mu-\lambda\cdot k)}{k!}$ (5.5)
For the EDN signal, $\mu=DCR^{*}\cdot t_{W}\cdot S$, where $DCR^{*}$ is equivalent DCR, $t_{W}$ is the average width of a single PE waveform, and S is the sensitive area of the SiPM in one channel. By replacing N in Equation 5.1 with GP in Equation 5.5 and fitting this function to the charge spectrum, the fitted value of $\lambda$ corresponds to the calibration result of IOCT.
The IOCT calibration result is shown in Figure 13, the bias is 1.40% and the standard deviation is 3.02%. According to the error propagation formula (Equation 5.2 to Equation 5.6), the bias caused by AP is 1.65%. The result indicates that the bias induced by multi-PEs EDN has been eliminated.
$Bias_{\rm APM,\lambda}=\frac{Bias_{\rm APM}}{1-P_{\rm IOCT}}$ (5.6)
Figure 12: Generalized Poisson Fit to the Charge Spectrum.
Figure 13: Bias Between IOCT Calibration Result and Mock Truth. The parameter calibrated for IOCT corresponds to the branch Poisson distribution parameter.
6 External Optical Crosstalk Calibration
The external optical crosstalk (EOCT) effect occurs when photons generated by avalanche electrons in APD escape from that APD and are detected by another APD in a different channel [Ref44, Ref30]. This paper presents a methodology for calibrating the EOCT rate and the distribution of EOCT emission angle in the TAO CD by switching the SiPM high voltage and using an LED external trigger to acquire data.
As shown in Equation 6.1, when all SiPMs in the TAO CD are operational, the electrical pulses generated within a SiPM include photoelectrons produced by the LED, DN, IOCT, AP, and EOCT from other SiPMs. Therefore, the charge $Q_{O}$ consists of contributions from the LED photon signal ($Q_{\rm LED}$), DN ($Q_{\rm DN}$), IOCT ($Q_{\rm IOCT}$), AP ($Q_{\rm AP}$) and EOCT ($Q_{\rm EOCT}$). When other SiPMs are turned off, the charge $Q_{C}$ only includes LED, DN, IOCT, and AP components. Thus, when acquiring "turn on" and "turn off" data over the same time duration, the difference between $Q_{O}$ and $Q_{C}$ corresponds to the contribution of EOCT. Here, the EOCT ratio is defined as $Q_{O}$ - $Q_{C}$ divided by $Q_{C}$.
$\begin{split}Q_{O}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}+Q_{\rm EOCT}\\ Q_{C}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}\\ Rate_{\rm EOCT}&=\frac{Q_{O}-Q_{C}}{Q_{C}}\end{split}$ (6.1)
In the TAO CD, the operating voltage of each SiPM is controlled via the high-voltage system. Specifically, an SiPM can be turned off by setting its voltage below the breakdown voltage. Since the number of active SiPMs in the CD varies before and after SiPM switching, data can not be acquired using the nhit self-trigger. Therefore, an LED external trigger system is employed. Since high voltage is applied to a single SiPM tile, the EOCT calibration in this study can only provide tile-level calibration results rather than channel-level results. (In TAO CD, each tile corresponds to two channels)
6.1 External Optical Crosstalk Rate Calibration
Three LED external trigger data sets were simulated to study the SiPM EOCT rate calibration. The first data set had all SiPMs turned on; the second had all SiPMs turned on but with the EOCT effect disabled in the simulation; the third had only one SiPM tile turned on. The EOCT rate calibration result can be derived from the first and third data sets, and the true EOCT rate in the simulation can be obtained from the first and second data sets. By comparing the calibration with the true value, the bias and standard deviation can be determined.
The EOCT rate calibration result is shown in Figure 14, the bias is 0.009% and the standard deviation is 0.096%.
Figure 14: EOCT Rate Calibration Result. The left figure presents the EOCT rate calibration value for each channel, with the definition of EOCT rate provided in Equation 6.1. The variation in EOCT rate across different layers is attributed to optical reflection at the SiPM surface. The middle and right figure shows the bias between the calibration value and the mock truth.
The bias and standard deviation are small because the calibrated EOCT rate (defined in Equation 6.1) is a directly observable quantity rather than a physical parameter such as the EOCT probability for each SiPM. The bias is quantified as the relative difference between the rates from Equation 6.1 and Equation 6.2 (In Equation 6.2, $Rate^{\rm truth}_{\rm EOCT}$ is the true EOCT ratio, $Q_{\rm EOCT}$ is the charge from EOCT, $Q_{C}$ is the charge measured when other SiPMs are turned off). Due to the LED sources can only be deployed along the central axis, the EOCT emission angle distribution must be calibrated to infer the response off the central axis.
$Rate^{\rm truth}_{\rm EOCT}=\frac{Q_{\rm EOCT}}{Q_{C}}$ (6.2)
6.2 External Optical Crosstalk Emission Angle Distribution Calibration
The external optical crosstalk emission angle is defined as the angle between the direction of EOCT photons escaping from a SiPM and the normal to the SiPM surface. In this study, the distribution of the SiPM EOCT emission angle is calibrated by using a reference SiPM to detect and quantify EOCT from other SiPMs positioned at different 5-degree angular intervals relative to the reference SiPM. The procedure is as follows: First, both the reference SiPM and the SiPMs within each 5-degree angular interval are turned on, and the total charge $Q_{O}$ on the reference SiPM is recorded. Next, only the reference SiPM is turned on, and data are acquired over the same duration to measure the total charge $Q_{C}$ on the reference SiPM. Finally, the EOCT response for each angular interval is obtained by subtracting $Q_{C}$ from $Q_{O}$.
The EOCT measurement response is the integrated response E from all SiPMs within each 5-degree angular interval. According to the mean value theorem for integral, the EOCT response $\overline{\epsilon(\bar{\theta})}$ in each 5-degree angular interval is given by Equation 6.3, where $\Delta\vec{S}$ is the area vector of each SiPM (with magnitude equals to the sensitive area and direction normal to the surface), $\vec{n}$ is the unit direction vector pointing from the emitting SiPM to the reference SiPM, $E(\theta)$ is the integrated EOCT response in each 5-degree angular interval, and $\epsilon(\theta)$ is the EOCT response at each angle.
$\begin{split}\overline{\epsilon(\bar{\theta})}&=\frac{E(\theta)}{\sum\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}\\ &=\frac{\sum\epsilon(\theta)\cdot\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}{\sum\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}\end{split}$ (6.3)
The reference SiPM selected for this study is located at the detector’s equator. This choice was made because apertures are present at the northern and southern poles of the TAO CD. To maximize the collection of EOCT photons, it is essential to position the reference SiPM away from these apertures. Therefore, the equatorial location was determined to be the optimal position.
The EOCT emission angle distribution calibration result is shown in Figure 15. The left part of Figure 15 presents the result when optical reflection on the SiPM surface is disabled in the simulation. In this case, the bias for most angular bins falls within the statistical uncertainties, except for the first bin. Since the SiPMs do not have a 100% fill factor and the EOCT emission angle distribution exhibits its most significant variation within the first bin, the average value over the sensitive area differs from that over the total area in this bin. The right part displays the result with optical reflection on the SiPM surface enabled. Here, discrepancies between the measured and true values are observed at large angles. However, as the EOCT photon distribution is predominantly concentrated at small angles, the impact of large-angle deviations on the experiment remains acceptable.
Figure 15: The EOCT Emission Angle Distribution Calibration Result. The two figures on the left display result without the SiPM surface reflection effect, while the two figures on the right show result including the reflection effect. The upper row uses a linear y-axis to visualize result in the small-angle range, and the lower row employs a logarithmic y-axis to present result in the large-angle range.
Since the LED light field distribution may be non-uniform, this study simulated such a non-uniform LED light field distribution, as illustrated in Figure 16. The corresponding EOCT emission angle distribution calibration result is shown in the left part of Figure 17. This result indicates that the non-uniform LED light field distribution affects the EOCT emission angle distribution calibration.
Figure 16: LED Non-Uniform Light Field Distribution Configuration in Simulation. The distribution follows a Gaussian distribution (Gaus(90,30)), where the parameters correspond to the polar angle values in the spherical coordinate system.
In the TAO calibration system, the $^{68}{\rm Ge}$ source produces uniform scintillation light in the LS, allowing us to correct the non-uniformity of LED light field. Data maps for both the $^{68}{\rm Ge}$ source and the LED source are acquired with all SiPMs turned on. During the EOCT emission angle distribution calibration, the charge measured by the reference SiPM in the "turn on" data is multiplied by a correction factor $\alpha$ defined in Equation 6.4. The summation is performed over all SiPMs within the corresponding angular range, where $Q_{{}^{68}{\rm Ge}}$ represents the charge in the $^{68}{\rm Ge}$ data map, and $Q_{\rm LED}$ represents the charge in the LED data map.
Figure 17: The EOCT Emission Angle Distribution Calibration Result by Using a Non-Uniform Light Field. The two figures on the left show result without correction, while the two figures on the right show the result with correction. The upper row uses a linear y-axis to visualize result in the small-angle range, and the lower row uses a logarithmic y-axis to display result in the large-angle range.
$\alpha=\frac{\sum Q_{{}^{68}{\rm Ge}}}{\sum Q_{\rm LED}}$ (6.4)
The corrected result is shown in the right part of Figure 17, where the influence of the non-uniform LED light field has been successfully corrected.
6.3 Correct Dark Count Rate Base on the External Optical Crosstalk Rate Calibration Result
The DCR calibrated in Section 4.1 includes contributions from EOCT; therefore, a correction is necessary. The correction method is provided by Equation 6.5, where $\beta$ is the correction factor, $\delta_{\rm EOCT}$ is the average charge response to EOCT, $\delta_{\rm IOCT}$ is the average charge response to IOCT, Q is the charge measured by the SiPM from each effect, $Rate_{\rm EOCT}$ is the calibrated EOCT ratio, and $\lambda$ is the branching Poisson constant for IOCT. The values of $\delta_{\rm IOCT}$ can be derived from the mean value of the generalized Poisson distribution [Ref31]. Because the charge contribution of AP in the TAO SiPM mass test result is very low [Ref32], its contribution is neglected. Since $Rate_{\rm EOCT}$ is calibrated at the tile level, both channels within the same tile use the same $Rate_{\rm EOCT}$ value.
$\begin{split}DCR&=DCR^{*}\cdot\beta\\ Rate_{\rm EOCT}&=\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}}\\ &\approx\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}}\\ &=\frac{\delta_{\rm EOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}{Q_{\rm LED}+Q_{\rm DN}+\delta_{\rm IOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}\\ \delta_{\rm EOCT}&=(1+\delta_{\rm IOCT})\cdot Rate_{\rm EOCT}\\ \delta_{\rm IOCT}&=\frac{\lambda}{1-\lambda}\\ \beta&=\frac{Q_{\rm DN}}{Q_{\rm DN}+Q_{\rm DN,\rm EOCT}}\\ &=\frac{Q_{\rm DN}}{Q_{\rm DN}+\delta_{\rm EOCT}\cdot Q_{\rm DN}}\\ &=\frac{1}{1+\delta_{\rm EOCT}}\\ &=\frac{1}{1+\frac{Rate_{\rm EOCT}}{1-\lambda}}\end{split}$ (6.5)
The result is shown in Figure 18. The bias is -0.40% and the standard deviation is 1.32%.
[…]
During the EOCT calibration process, some data is acquired with only a single SiPM turned on. This data can be used to calibrate the DCR without the influence of the EOCT effect. It provides a reference for independent verification and comparative analysis.
7 Conclusion & Limitation
7.1 Conclusion
This study presents a systematic investigation of the channel-level and tile-level calibration of key parameters for the silicon photomultiplier (SiPM) and its associated electronics in the TAO central detector. The data sets used in this study are summarized in Table 2.
At the channel level, the calibration approaches for individual parameters are as follows:
  1. 1.
    Dark Count Rate (DCR): Calibrated by counting the number of hits recorded before the trigger time.
  2. 2.
    Time Offset: Calibrated using external trigger LED events with the source positioned at the center of the LS. The time offset for each channel is determined by analyzing the difference in the most probable hit time value for each channel.
  3. 3.
    Relative Photon Detection Efficiency (PDE): Calibrated using $^{68}{\rm Ge}$ events with the source positioned at the center of the liquid scintillator (LS), where the relative PDE between channels is determined by the relative number of zero-hit events across these channels.
  4. 4.
    Gain: The charge spectrum is analyzed using a multi-Gaussian function fit. The gain value is determined by the separation between adjacent Gaussian peaks in the multi-Gaussian function.
  5. 5.
    Internal Optical Crosstalk (IOCT): An initial approach calibrated the IOCT by analyzing the proportion of multi-photoelectrons (multi-PEs) components in DN events. However, this method is susceptible to inaccuracies due to intrinsic multi-PEs DN. To overcome this limitation, an improved method was developed that employs a generalized Poisson distribution fit to the DN charge spectrum, resulting in a more reliable IOCT calibration.
At the tile level, two critical parameters are calibrated:
  1. 1.
    External Optical Crosstalk (EOCT) Rate: The operating state of SiPMs is controlled by using the high-voltage system. The EOCT rate for a target SiPM is determined by measuring the difference in response when all other SiPMs are turned on and turned off.
  2. 2.
    External Optical Crosstalk Emission Angle Distribution: Calibration is performed by selectively turning on/off SiPMs within a specific angular range relative to a target SiPM. The bias is less than 4% within the main angular range of EOCT emission photons.
    The bias and standard deviation of all calibration parameter results are summarized in Table 3.
Table 3: The Bias and Standard Deviation of Each Calibrated Parameter
Calibrated Parameter Bias Standard Deviation
DCR (Uncorrected) 23.6% 1.66%
DCR (Corrected) -0.40% 1.32%
Gain 0.084% 0.094%
Time Offset 0.027 ns 0.031ns
Relative PDE <3% 111between the channel at pole and equator. 0.17%
IOCT(Multi-PEs) 5.70% 2.73%
IOCT(GP Fitting) 1.40% 3.02%
EOCT Rate 0.009% 0.096%
7.2 Limitation
This paper does not present a calibration method for the AP. As shown in Figure 19, AP contributes a fine structure to the tail of each PE peak in the charge spectrum. However, electronic noise and the charge reconstruction algorithm implemented on the FPGA can produce similar fine structures. Since the noise level and charge reconstruction algorithm differ between simulated and real data, and considering that the AP rate measured in the TAO SiPM mass test is very low [Ref32], this paper does not provide a calibration method for AP. A dedicated calibration method will be proposed in future work based on real data.
Figure 19: Charge Spectrum Distributions with 1% AP and without AP. The red line represents the case with AP, while the blue line represents the case without AP. It can be observed that AP contributes a fine structure to the tail of each PE peak in the charge spectrum.
Similarly, due to the electronic environment, the charge and time reconstruction algorithm in reality differ from those used in the simulation. The calibration method presented in this paper will be optimized based on real data in future work, and the corresponding calibration results will be provided subsequently.
This paper does not provide a method for correcting the optical reflection effect on the SiPM surface. A solution will be updated in future work based on Geant4 detector optical simulation.
Acknowledgments
We are gratefully for the JUNO-TAO and the technical staff of the participating institutions. This work was supported by National Natural Science Foundation of China under Grant Number 12275281.