### v2

# Channel-Level Calibration Methods of Silicon Photomultiplier for JUNO-TAO Central Detector

Abstract
The Taishan Antineutrino Observatory (TAO or JUNO-TAO) is a satellite observatory for the Jiangmen Underground Neutrino Observatory (JUNO), located 44 meters away from the No.1 reactor of the Taishan Nuclear Power Plant. TAO can measure the reactor antineutrino energy spectrum with excellent energy resolution (better than 2% at 1 MeV) using state-of-the-art Silicon Photomultipliers (SiPMs) operated at low temperature. To achieve this goal, the SiPMs (together with their readout electronics) must be well calibrated. This paper presents the channel-level calibration methods for the dark count rate (DCR), relative photon detection efficiency (PDE), time offset, gain, and internal optical crosstalk (IOCT) of the SiPMs based on charge and time information of the collected events. For the calibration of the external optical crosstalk (EOCT), in terms of its rate and emission angle distribution, a novel method is proposed by switching on and off different groups of SiPMs with an LED placed in the detector. Using one million simulated events, the expected calibration biases are evaluated for all the aforementioned parameters: relative PDE (2.1%), IOCT (1.4%), DCR (0.4%), EOCT Rate ($0.1\%$), gain ($0.1\%$), time offset (<0.2 ns). The emission angle distribution of the EOCT photons could be measured with a bias of less than 4% in main angular range.
1 Introduction
Before the precise measurements of the reactor neutrino energy spectrum by experiments such as Daya Bay [6], RENO [11], and Double Chooz [1], the primary theoretical models for calculating the expected reactor neutrino energy spectrum were the Summation model [14] based on nuclear database data [17, 5, 36, 32, 40, 31] and the $\beta$ conversion method such as Huber-Muller model [22, 30] based on the ILL High Flux Reactor measurement data [38, 37, 19]. However, according to the measurement results from the Daya Bay, RENO, and Double Chooz experiments, both the Summation model and the Huber-Muller model exhibit systematic deviations from the experimental results in terms of the predicted energy spectrum flux and the shape of the energy spectrum at 5 MeV. Although Estienne and Fallot have improved the Summation model [13]; and Kopeikin, Skorokhvatov, Titov [23]; Hayen, Kostensalo, Severijns, Suhonen [21]; Giunti, Li, Ternes, Xin [16] have refined the Huber-Muller model, these models currently can not simultaneously explain the systematic deviations in both flux and spectrum shape. Moreover, none of models can reliably predict the fine structures in the reactor antineutrino spectrum [12]. If the JUNO experiment directly uses these predicted spectra as inputs, it will introduce significant uncertainties in the energy spectrum, which will affect the sensitivity of the neutrino mass ordering measurement [3]. To reduce the model dependence of the reactor neutrino energy spectrum, JUNO has deployed a satellite detector, the Taishan Antineutrino Observatory (TAO or JUNO-TAO) at a baseline of 44 meters away from Unit 1 of the Taishan Nuclear Power Plant [2].
[…]
To achieve the aforementioned physical goals, it is crucial to attain the designed energy resolution for the TAO experiment. However, the negative effects of SiPMs can directly impact the energy resolution. For instance, according to the TAO Conceptual Design Report, dark noise and optical crosstalk (OCT) contribute to a degradation of 0.75% and 0.5% on the energy resolution at 1 MeV [2]. In addition to affecting the energy resolution, these two effects can cause deviations of 1 mm and 2 mm on the vertex reconstruction, as reported by Shi [39]. Therefore, accurate calibration of all relevant SiPM parameters for each readout channel is of crucial importance.
[…]
This paper investigates and presents channel-level calibration methods for SiPMs installed in the TAO CD, including calibration of the DCR, relative PDE, internal optical crosstalk (IOCT) rate, gain, and time offset. Additionally, a method based on selectively turning on and off different groups of SiPMs utilizing the fine control capability of the SiPM high-voltage system and an LED calibration source is proposed to calibrate the EOCT rate and emission angle distribution. This paper is conducted using simulated data generated by the TAO offline software.
This paper is structured as follows: Section 2 introduces the TAO detector; Section 3 presents the simulated model; Section 4 describes the calibration of the DCR, relative PDE, and time offset based on hit time information; Section 5 details the calibration of gain and IOCT rate based on charge information; Section 6 introduces a novel method for EOCT rate and emission angle distribution calibration; Section 7 discusses the SiPM parameter uncertainties induced by temperature fluctuation and the impact of uncertainties originating from both temperature fluctuation and the calibration medthon on the reconstructed vertex and energy; Section 8 summarizes the research work presented in this paper.
2 TAO Detector
The TAO detector structure is shown in Figure 1. It consists of three sub-detectors: the central detector, water tanks and the top veto tracker.
Figure 1: Design of the TAO detector.
[…]
Surrounding the central detector are three water tanks. Each tank is filled with 50 tonnes of pure water, which can shields external radioactive background and produces water Cherenkov light signals when muons pass through. The tanks are equipped with 300 3-inch PMTs, which detect the water Cherenkov signals to veto muons.
Above the central detector is the top veto tracker [27, 28], which consists of four layers of plastic scintillators and is designed for muon veto.
3 TAO Central Detector Simulated Model
This paper adopts simulated data produced with the TAO offline software to investigate calibration methods for SiPM parameters, and this section details the corresponding simulation model.
The TAO offline software is a dedicated framework for experimental data processing, physics data analysis, and simulated data production for the TAO experiment. As a core offline component of the experiment, it is built on the SNiPER framework [26]. For the simulated data production in this work, three core modules are employed: physics event generation, detector simulation, and electronics simulation.
3.1 Physics Event Generator
The first step of the full Monte Carlo (MC) simulation chain for the TAO is to produce primary particles using the physics event generator. The generator used in the TAO is directly inherited from the well-validated offline framework of the JUNO experiment. As documented in the JUNO simulation paper [25], the JUNO framework supports three primary types of physics event generators: the particle gun, HepEvt, and GENIE.
In this work, we adopt the particle gun, a dedicated tool that generates particles with well-defined species, momenta, and positions. We generate one million particle events for each SiPM parameter, and the particle events information corresponding to each SiPM parameter is summarized in Table 1.
Table 1: Summary of the simulated data used for the channel-level calibration of the SiPM parameters
SiPM Parameter Particle Position Trigger Type
Dark Count Rate Arbitrary1 Arbitrary Self Trigger
Gain Arbitrary Arbitrary Self Trigger
Time Offset LED CD Centre External Trigger
Relative PDE $^{68}\text{Ge}$ CD Centre Self Trigger
Internal Crosstalk Arbitrary Arbitrary Self Trigger
External Crosstalk LED CD Centre External Trigger
  • 1
    We use electrons uniformly distributed within the liquid scintillator with energies following a uniform distribution in the range of 1–10 MeV to represent arbitrary events.
3.2 Detector Simulation
Following the physics event generation, detector simulation is performed as the next step of the simulation chain. The detector simulation module in the TAO offline software is built based on the Geant4 (version:10.04.p02) [4] toolkit. Firstly, the TAO detector geometry is constructed in Geant4 according to the detector structure of the TAO shown in Figure 1. Subsequently, the physics processes of particles propagating inside the detector are defined. All physics processes are consistent with those implemented in the detector simulation of the JUNO [25], except for two dedicated optical processes: the scintillation light emission induced by energy deposition of particles in the gadolinium-doped liquid scintillator, and the physical process of optical photon detection by SiPMs.
For the simulation of optical processes in the liquid scintillator, a quenching effect occurs when charged particles deposit energy in the liquid scintillator. This quenching effect is described by Birks’ law [7], and the Birks constant of the liquid scintillator adopted in the TAO is determined from dedicated experimental measurements [47]. The remaining deposited energy, after excluding the quenched component, drives the scintillation light emission of the liquid scintillator. The wavelength distribution of the emitted scintillation photons is determined by the 2,5-diphenyloxazole (PPO) and p-bis-(o-methylstyryl)-benzene (bis-MSB), and the wavelength spectrum implemented in the TAO detector simulation is fixed based on the experimentally measured in Ref. [49]. The optical attenuation length of the TAO liquid scintillator is jointly determined by the linear alkylbenzene, PPO and bis-MSB, and the value adopted in the TAO simulation is experimentally determined as specified in Ref. [49]. In addition, the light yield and scintillation time profile of the TAO liquid scintillator are experimentally characterized and determined as described in Ref. [44].
For the simulation of optical processes for SiPMs, when optical photons hit on the surface of a SiPM, they can be detected or reflected by the SiPM. The former is determined by the PDE of the SiPM, while the latter is governed by the optical reflectivity of the SiPM. Both parameters implemented in the TAO detector simulation are derived from dedicated experimental measurements: the PDE is taken from Ref. [45], and the optical reflectivity is adopted from Ref. [43].
Finally, the SiPMs detect both the light produced by energy deposition of particles in the liquid scintillator and the Cherenkov light generated via Cherenkov radiation, and store the corresponding detection data as SiPM hit information.
3.3 Electronics Simulation
Following the generation of SiPM hit information via detector simulation, electronics simulation is performed as the subsequent stage of the full simulation chain. The TAO electronics simulation workflow is illustrated in Figure 2, which takes the SiPM hit information produced by detector simulation as its input. The hit information is fed into the TaoPdSimAlg algorithm, which converts the SiPM hit information from each detector simulation event into electrical pulse data. Each electrical pulse is characterized by three core parameters: relative amplitude, timestamp, and type identifier. The relative amplitude is determined by the single photoelectron (PE) gain and single PE charge resolution of the SiPM. For the timestamp, the event-level timestamp of each electronics simulation event is obtained by sampling based on the physics event rate; the timestamp of each electrical pulse is then calculated by summing the event-level timestamp, the corresponding SiPM hit time, and the contributions from the electronics time offset and time smearing effects. The type identifier is used to indicate the specific SiPM channel or electronics process from which the electrical pulse originates. Thereafter, additional electrical pulses induced by the intrinsic SiPM effects, including dark noise, OCT, and afterpulse (AP), are generated based on the corresponding dedicated models. Details of the SiPM effect models are elaborated in Section 3.4.
Figure 2: TAO Electronics Simulation Workflow.
Following the generation of all electrical pulses, trigger decision and event packaging are performed within the TaoTrigSimAlg algorithm. The TAO electronic system supports two distinct trigger modes: self-trigger mode and external trigger mode. For the self-trigger mode, trigger decision is implemented based on a 300 ns trigger time window. This time window slides along the time axis, when the number of channels in which electrical pulses are detected within the this time window exceeds the trigger threshold (set to 1000 in the TAO electronics simulation), a valid trigger will be generated, and the start time of the trigger time window is defined as the trigger time. For the external trigger mode, the internal threshold-based trigger decision logic is bypassed, and a forced trigger is generated for each individual detector simulation event. The timestamp sampled from the corresponding detector simulation event is directly adopted as the electronics trigger time. For both trigger modes, a fixed readout window spanning from 100 ns before the trigger time to 900 ns after the trigger time is applied, and all electrical pulses falling within this window are packaged into a single electronic event.
Following the generation of electronics event, the event is fed into the TaoElecSimAlg algorithm for waveform simulation, which converts each electrical pulse into an electrical pulse waveform. The amplitude of the waveform is determined by the relative amplitude of the electrical pulse, and white noise with a signal-to-noise ratio of 10 is added to the waveform. Subsequently, SiPM waveform identification and charge-time calculation are performed. A valid SiPM waveform is identified when three consecutive sampling points exceed the preset threshold (set at 50% of the single PE waveform amplitude). The charge corresponding to the waveform is obtained via waveform integration, and the timing of the waveform is determined by linear fitting of the waveform rising edge, where the intersection of the fitted rising edge and the baseline is defined as the waveform timing. After the charge-time calculation is completed, three parameters of each waveform are obtained: the charge, the time relative to the trigger time, and the associated electronics channel number. Finally, the SiPM parameters of each channel are calibrated based on the timing and charge information.
In the TAO simulation, the true values of each SiPM parameter are known. Accordingly, we quantify the calibration performance using two core metrics: the calibration bias and the standard deviation. The calibration bias is defined as the mean difference between the calibrated value and the corresponding simulated truth, averaged over all channels. The calibration standard deviation is calculated as the standard deviation of the residuals (calibrated value minus the corresponding simulated truth) across all channels.
3.4 SiPM Simulation Modeling
3.4.1 Dark Noise
Dark noise refers to the effect in which a SiPM spontaneously generates avalanche current in the absence of photon hits. It arises from two distinct physical processes: one is thermal current, which is generated when valence band electrons in the SiPM are excited into the conduction band by thermal energy, and trigger an electron avalanche to produce avalanche current during this process; the other is tunneling current, which occurs when valence band electrons in the SiPM have a probability of tunneling into the conduction band via the quantum tunneling effect under the action of an electric field, and subsequently induce avalanche current [33].
[…]
3.4.2 Time Offset
In the TAO experiment, differences in the length of signal cables for SiPMs at different positions, and potential time asynchrony between the ADCs used in the TAO readout system. These effects lead to relative time offsets between different electronics channels. In the TAO electronics simulation, the time offset is implemented by adding distinct time offset constants to the timestamps of electrical pulses corresponding to different channels. In this study, the time offset is configured to 0 ns for the upper hemisphere channels and 20 ns for the lower hemisphere channels.
3.4.3 Photon Detection Efficiency
The SiPM PDE is defined as the ratio of the number of photons detected by the SiPM to the number of photons incident on the SiPM surface. The SiPM PDE is generally characterized by three factors [20, 48]:
$PDE(\lambda,V)=QE(\lambda)\cdot Prob_{\rm GM}(V)\cdot FF$ (3.1)
where QE is the Quantum Efficiency, which varies with the wavelength of incident light; $Prob_{\rm GM}$ is the probability of triggering an electron avalanche effect, which is related to the internal electric field strength of the SiPM and thus to the operating voltage of the SiPM; and FF (Fill Factor) is a geometric factor used to evaluate the impact of the SiPM geometry on the photon detection efficiency. For the TAO simulation, the detector geometry is constructed from engineering drawings, and the geometric effect is automatically set by Geant4.
The PDE used in the TAO detector simulation is set to 50%, which is generally consistent with the measured results from TAO SiPM mass test [42]. This PDE value is configured on the optical surface of the Geant4 sensitive detector corresponding to the SiPMs in the TAO detector simulation.
3.4.4 Gain
Gain is defined as the total charge of the waveform, which is generated after a photon hits the SiPM, is detected, produces an avalanche current, and is amplified and shaped by the electronics system.
Due to various stochastic processes in the SiPM and the electronics, there is a fluctuation in the amplification factor that converts a single photoelectron into a single photoelectron waveform, which is the charge resolution of the SiPM and the electronics. In the TAO electronics simulation, a 15% Gaussian smearing is added to the relative amplitude of the electrical pulse to represent the 15% charge resolution of the TAO SiPM and electronics reported in the TAO Conceptual Design Report [2].
3.4.5 Optical Crosstalk
OCT of the SiPM arises from the fact that avalanche electrons have a finite probability of emitting photons during acceleration in the electric field of the SiPM. These photons have a certain probability of leaving the Avalanche Photodiode (APD) where they are generated, impinging on other APDs, and triggering new avalanche signals [29, 35].
If the APD hit by the OCT photon and the APD that generates the photon belong to the same electronics channel, this effect is defined as IOCT. If the APD hit by the OCT photon and the APD that generates the photon belong to different electronics channels, this effect is defined as EOCT.
According to the study by Vinogradov [41], the number of OCT events generated by a signal follows a cascaded Poisson distribution. For a single avalanche signal, the total number of the primary signal and the OCT events it induces follows a Borel distribution:
[…]
where $k$ represents the total number of SiPM electrical pulses induced by one primary avalanche signal and all subsequent OCT events triggered by the primary signal, and $\lambda_{\rm OCT}$ is the Poisson parameter characterizing the intrinsic OCT probability per avalanche event.
The physical meaning behind this mathematical formulation can be intuitively interpreted as a multi-generation cascaded avalanche process: a primary avalanche signal produces a set of first-generation OCT events through a Poisson process with rate parameter $\lambda_{\rm OCT}$. Each of these first-generation OCT events will trigger an independent avalanche, which can further generate second-generation OCT events via an identical Poisson process with the same rate parameter $\lambda_{\rm OCT}$. This iterative cascade process continues until no additional OCT events are generated in a given generation. Mathematically, the total count of the primary signal and all cascaded OCT events across all generations converges to the Borel distribution given above.
More generally, for avalanche signals generated by a Poisson process with a Poisson coefficient of $\mu$, the sum of the number of avalanche signals and the number of OCT events generated by the avalanche signals mathematically follows a generalized Poisson distribution:
$GP(k,\mu,\lambda)=\frac{\mu\cdot(\mu+\lambda\cdot k)^{k-1}\cdot exp(-\mu-\lambda\cdot k)}{k!}$ (3.3)
In the TAO electronics simulation, IOCT and EOCT are simulated separately. For IOCT, the corresponding number of IOCT pulses is generated via Borel distribution sampling. IOCT will be induced by every electrical pulse generated from a SiPM photon hit, a dark noise electrical pulse, and an EOCT electrical pulse. According to the study by Guan et al. [18], the IOCT of HPK SiPMs in linear alkylbenzene is very low, with a Poisson coefficient $\lambda_{\rm IOCT}$ less than 0.1%, which makes this effect negligible and no corresponding calibration required. The TAO SiPM batch tests measured the in-air IOCT probability of the SiPMs, which value is approximately 15%. We directly adopted this value in the TAO electronics simulation, to investigate how to calibrate this parameter under the scenario where the IOCT effect is non-negligible.
The simulation of EOCT is implemented by performing Poisson sampling on each generation of electrical pulses. Each electrical pulse, including those induced by SiPM photon hits, dark noise, IOCT, is subjected to Poisson sampling with a coefficient of $\lambda_{\rm EOCT}$ to generate first-generation EOCT pulses. These newly generated EOCT pulses, together with the IOCT produced by them, are further processed with Poisson sampling of the same coefficient $\lambda_{\rm EOCT}$ to produce a new generation of EOCT pulses. Such a cascaded process proceeds iteratively until no additional EOCT pulses are generated. In the TAO electronics simulation, the $\lambda_{\rm EOCT}$ is adopted from the measurement of EOCT for HPK SiPMs reported in the study by Guan et al. [18], with a value of 0.2447.
The electronics channel ID hit by EOCT is obtained by simulating a specific photon emission distribution on the SiPM surface and observing the photon hit responses of other SiPMs. Taking the SiPM with SiPM ID = 0 as an example, one million optical photons are simulated using a particle gun on the surface of the corresponding SiPM. The angle between the photon propagation direction and the surface normal is sampled from a Gaussian distribution with a mean of $0^{\circ}$ and a standard deviation of $30^{\circ}$. If the sampled angle larger than $90^{\circ}$, the sampling is repeated until angle less than or equal to $90^{\circ}$.
Following the above simulation procedure, the distribution of SiPM ID hit by EOCT generated from the SiPM with ID = 0 is shown in Figure 3. Since the photons generated by EOCT are emitted predominantly along the normal direction of the SiPM surface, the SiPM facing the EOCT-emitting SiPM within each layer of the SiPMs receive the largest number of EOCT photons. This results in a multi-peak structure in the SiPM ID distribution, where each peak corresponds to the EOCT response received by the SiPMs in each layer. The electronics channel ID hit by EOCT is then obtained by sampling from this histogram.
Figure 3: Hit SiPM ID Distribution from EOCT Hits Generated by the SiPM ID = 0 Channel.
It is worth noting that since the coverage of TAO SiPMs is not 100%, the total number of photon hits is less than one million. The ratio of the total number of photon hits to one million is defined as the geometric factor for EOCT. The actual Poisson coefficient for EOCT is 0.2447 multiplied by the geometric factor corresponding to the SiPM. The geometric factor varies for SiPMs at different positions. The distribution of the geometric factor for all 4024 SiPMs is shown in Figure 4, the mean value is 0.7429.
Figure 4: Geometric Factor Distribution for EOCT.
3.4.6 Afterpulse
During the avalanche discharge process of the APD, there is a probability that electron-hole pairs are trapped by lattice defects. The trapped carriers will be re-released, which in turn triggers a new avalanche process; this phenomenon is AP [35]. The AP in SiPMs can be divided into two categories. One is delayed AP, which has a long time interval from the main avalanche signal (> 1 $\mu\text{s}$). For the TAO SiPM, the measured delayed AP is negligible [45]. The other is non-delayed AP. Due to the short time interval between this type of AP and the main signal, the APD is still in the avalanche recovery phase at this time, so the gain of the AP is lower than that of the primary avalanche pulse.
Based on the result of the TAO SiPM mass test [42], the AP probability of the TAO SiPM is very low (< 1%). In the TAO simulation, we have implemented a 1% AP effect. The time interval $\Delta t$ between the AP and the primary pulse, as well as the relative gain of the AP $\text{Gain}_{AP}$ with respect to the primary pulse, both follow an exponential distribution [10], as shown in Equation 3.4 and Equation 3.5, respectively.
$f(\Delta t)=e^{-\frac{\Delta t}{\tau_{AP}}}$ (3.4)
$\text{Gain}_{AP}=1-e^{-\frac{\Delta t}{\tau_{rec}}}$ (3.5)
where $\tau_{\text{AP}}$ is the AP time constant, and $\tau_{\text{rec}}$ is the avalanche recovery time constant. Since these two time constants were not measured in the TAO SiPM mass test, we adopted the values reported in Ref. [34], which are 20 ns and 13 ns, respectively.
4 Calibration Based on Hit Time Information
4.1 Dark Count Rate
The DCR can be estimated by counting the recorded hits in the plateau region before the main signal hits from physics events. This plateau region corresponds to the pre-trigger time, which is the area to the left of the red line (indicating the trigger time) in Figure 5. The DCR can be calculated using Equation 4.1:
$\text{DCR}=\frac{n}{N\cdot t\cdot S}$ (4.1)
Where $n$ denoted as the total number of the counted hits in the plateau, $N$ as the number of physics events, $t$ as the duration of the plateau before the trigger timestamp (from -100 ns to 0 ns), and $S$ as the sensitive area of the SiPM.
Figure 5: Hit time distribution from self-triggered events. The trigger time corresponds to 0 ns, which is indicated by the red line in the figure. The hits from physics events are concentrated around 300 ns, as determined by the 300 ns width of the trigger time window. The plateau preceding the hits of physics events corresponds to the dark noise plateau.
However, the hits observed in the plateau range may also include minor contributions from hits generated by the EOCT of dark noise from other SiPMs, as well as hits from delayed APs. For the TAO SiPM, the delayed AP has been measured to be very weak [45]. Therefore, only EOCT contamination needs to be considered in the DCR calibration.
The DCR calibration result is shown in the left part of Figure 6. Since the SiPMs are mounted in circle around the central $z$-axis of the detector and layer by layer along the central $z$-axis, the channels have azimuth number and layer number as its coordinate. A significant bias of 23.6% with the standard deviation of 1.7%. However, when the EOCT effect is disabled, the bias reduces is substantially reduced to -0.2% as presented in the right part of Figure 6. This clearly demonstrates that the EOCT effect introduces a considerable bias in the calibration result. A correction method will be presented in Section 6.3.
Figure 6: DCR Calibration Results. The left two panels: calibration results with EOCT included. The right two panels: calibration results with EOCT excluded. The upper two panels: comparison between the calibration results and the simulation truth in each channel. The lower two panels: 1D histogram filled with each bin value from the upper 2D histogram.
4.2 Time Offset
The hit time of the SiPM obtained via the electronic readout satisfies Equation 4.2, where $t_{R}$ is the hit time from the electronic readout. It is calculated as $t_{R}=t_{A}-t_{T}$, where $t_{A}$ is the absolute timestamp and $t_{T}$ is the trigger timestamp. The absolute timestamp consists of several components: $t_{P}$ is the time when the particle reaches and deposits energy in the liquid scintillator; $t_{\rm LS}$ is the time when the liquid scintillator emits scintillation photons; $t_{\rm tof}$ is the time of flight of scintillation photons to the SiPM; the electronic processing time, which is divided into two parts: $t_{\rm same}$ and $t_{\rm timeoffset}$. Here, $t_{\rm same}$ represents the identical portion of the processing time across all electronic channels, while the differing portions constitute the time offset ($t_{\rm timeoffset}$) in different channels.
$\begin{split}t_{R}&=t_{A}-t_{T}\\ t_{A}&=t_{P}+t_{\rm LS}+t_{\rm tof}+t_{\rm same}+t_{\rm timeoffset}\\ t_{\rm timeoffset}&=t_{R}+t_{T}-t_{P}-t_{\rm LS}-t_{\rm tof}-t_{\rm same}\end{split}$ (4.2)
Since the TAO experiment only requires correcting the relative time offsets between electronic channels [2] to ensure that the timestamps of all channels are synchronized, this study deploys an LED source at the center of the liquid scintillator to calibrate the relative time offset between different channels. The LED source acquires events using an external trigger method, ensuring that the trigger time $t_{T}$ is identical across all channels. Therefore, $t_{T}$ does not affect the calibration of relative time offset. The $t_{P}$ and $t_{\rm LS}$ are statistically identical across all channels, so the $t_{P}$ and $t_{\rm LS}$ does not influence relative time offset calibration. Because the LED is deployed at the center of the liquid scintillator, the distance from the liquid scintillator center to each SiPM channel is the same, meaning the $t_{\rm tof}$ is identical across all channels. Consequently, the relative time offsets between different channels can be derived from the relative $t_{R}$ differences for each channel.
As shown in Figure 7, we use the double-sided Crystal Ball function to fit the distribution of $t_{R}$. The mean value of the fitted function corresponds to the $t_{R}$ value for the channel. When a reference channel is selected, the time offset of $\rm channel_{\rm i}$ relative to the reference channel ($t^{\rm chid=i}_{\rm timeoffset}$) is the readout time of $\rm channel_{\rm i}$ ($t^{\rm chid=i}_{R}$) minus the readout time of the reference channel ($t^{\rm RefCh}_{R}$).
$t^{\rm chid=i}_{\rm timeoffset}=t^{\rm chid=i}_{R}-t^{\rm RefCh}_{R}$ (4.3)
Figure 7: Double-Sided Crystal Ball Function Fit to the Time Spectrum. The mean value of the fitted function corresponds to the $t_{R}$ value.
The time offset calibration result is shown in Figure 8, the bias is 0.027 ns and the standard deviation is 0.031 ns.
Figure 8: Time Offset Calibration Results. The upper panel: channel-level calibration results. The middle panel: comparison between the calibration results and the simulation truth in each channel. Lower panel: 1D histogram filled with the each bin value from the middle 2D histogram.
In the actual TAO CD, the the calibration source’s position may be offset from its deployed position, and this offset contributes to a bias in the time of flight. Specifically, a 1 cm offset results in a 0.13 ns bias between the SiPMs located at the south and north poles of the CD. Considering that the time resolution of the TAO electronics exceeds 1 ns [2], both the calibration bias and the calibration source’s positional offset from the liquid scintillator center are acceptable.
4.3 Relative Photon Detection Efficiency
As shown in Equation 4.4, the PDE of a SiPM is defined as the ratio of the number of photons detected by the SiPM ($N_{\rm Det}$) to the number of photons incident on the SiPM surface ($N_{\text{photon}}$):
$\text{PDE}=\frac{N_{\rm Det}}{N_{\text{photon}}}$ (4.4)
Since the number of photons incident on the SiPM surface depends on the particle type, particle energy, light yield of the liquid scintillator, and light attenuation length of the liquid scintillator, calibrating the absolute PDE of each SiPM is impractical. However, calibrating the relative PDE is feasible. The TAO calibration system employs three types of calibration sources: an LED source, a $^{68}{\rm Ge}$ source and a combined $\gamma$ source [46]. The light field produced by LED source may be not isotropic, and the combined $\gamma$ source consists of multiple $\gamma$ sources with different energies. In contrast, the $^{68}{\rm Ge}$ source is a positron annihilation $\gamma$ source. Due to the presence of the source shell, positrons annihilate inside the shell, producing a pair of $\gamma$ emitted in opposite directions, which enhances the uniformity of the light field. Therefore, the $^{68}{\rm Ge}$ source is used for relative PDE calibration.
Deploying the $^{68}{\rm Ge}$ source at the center of the liquid scintillator, the number of photons detected by each channel follows a Poisson distribution. The mean number of photons detected by the SiPM is the parameter of the Poisson distribution, which satisfies:
$N_{\rm Det}=-ln(\frac{N_{\text{PE}=0}}{N_{\text{Event}}})$ (4.5)
Where $N_{\text{PE}=0}$ is the number of events without $^{68}{\rm Ge}$ photon hits in the channel, and $N_{\text{Event}}$ is the number of events. Since the $^{68}{\rm Ge}$ source is deployed at the center of the liquid scintillator and the light emission from the liquid scintillator is isotropic, the average number of incident photons on the surface of each channel is the same. According to Equation 4.4, the relative PDE of each channel corresponds to the relative $N_{Det}$ value.
However, the dark noise effect can also generate hits. If the DCR varies across different readout channels, the number of hits contributed by dark noise will differ between channels, which in turn leads to variations in $N_{\text{PE}=0}$, thereby introducing a bias in the calibration results of the relative PDE. For this reason, a dedicated correction for the dark noise effect is required.
The hit time distribution is shown in Figure 5. After applying the time offset correction, we count hits from 0 ns to 160 ns, which are purely contributed by dark noise. Then, also count the hits from 240 ns to 400 ns, which include contributions from both dark noise and $^{68}{\rm Ge}$. The probability of a dark noise occurring within the 160 ns time window, denoted as $P_{\rm DN}$, is defined as the ratio of the number of events containing dark noise $N_{\rm DN}$ to the $N_{\text{Event}}$. The probability of observing either dark noise or $^{68}{\rm Ge}$ (or both), denoted as $P_{\rm Ge,DN}$, is defined as the ratio of the number of events containing dark noise, or $^{68}{\rm Ge}$, or both occurring $N_{\rm Ge,DN}$ to the $N_{\text{Event}}$.
[…]
Dark noise and $^{68}{\rm Ge}$ are statistically independent. Therefore, the probability of the union of these two events $P_{\rm Ge,DN}$ equals the sum of the probabilities of three mutually exclusive scenarios: dark noise occurs without $^{68}{\rm Ge}$, $^{68}{\rm Ge}$ occurs without dark noise, and both dark noise and $^{68}{\rm Ge}$ occur simultaneously.
$\begin{split}P_{\rm Ge,DN}=P_{\rm DN}\cdot(1-P_{\rm Ge})+P_{\rm Ge}\cdot(1-P_{\rm DN})+P_{\rm Ge}\cdot P_{\rm DN}\end{split}$ (4.7)
Meanwhile, the probability of no $^{68}{\rm Ge}$ occurring within the 240 ns to 400 ns time window, denoted as $1-P_{\rm Ge}$, is defined as the ratio of the $N_{\text{PE}=0}$ to the $N_{\text{Event}}$.
$\begin{split}N_{\text{PE}=0}&=(1-P_{\rm Ge})\cdot N_{\text{Event}}\end{split}$ (4.8)
where $P_{\rm Ge}$ is derived from Equation 4.6 and Equation 4.7, and is given by:
$\begin{split}P_{\rm Ge}&=\frac{P_{\rm Ge,DN}-P_{\rm DN}}{1-P_{\rm DN}}\end{split}$ (4.9)
In this simulation, we set four different DCR values: 20, 40, 60, and 80 ${\mathrm{mm}}^{2}$. The results before and after applying this correction method are shown in Figure 9, this method can effectively correct the bias caused by different DCR levels.
Figure 9: Relative PDE Calibration Results Before and After DCR Correction. The left and right panels present the comparison between the PDE calibration results and the simulation truth before and after DCR correction, respectively.
Figure 10: Relative PDE Calibration Results Before and After Accounting for the Optical Reflection Effect of the SiPM. The left two panels: calibration results with the SiPM optical reflection effect included. The right two panels: calibration results with the SiPM optical reflection effect excluded. The upper two panels: comparison between the calibration results and the simulation truth in each channel. The lower two panels: 1D histogram filled with each bin value from the upper 2D histogram.
After Correcting the dark noise differences in each channel, the relative PDE calibration result is shown in the left part of Figure 10. There is an approximate 3% bias between the channels at the pole and equator, which is caused by optical reflection on the SiPM surface. As shown in the right part of Figure 10, this bias disappears after disabling the optical reflection on the SiPM surface in the detector simulation. This bias will be corrected based on the result of the detector optical model and simulation in the future.
5 Calibration Based on Charge Information
5.1 Gain
The charge generated by a single PE waveform follows a Gaussian distribution, and the individual photons detected by the APDs in the SiPM are independent. Consequently, the total charge resulting from the simultaneous detection of multiple PEs also follows a Gaussian distribution, and the overall distribution is a multi-Gaussian distribution (Equation 5.1 where f is the multi-Gaussian function, $N_{i}$ is the number of hits corresponding to each PE, $\mu_{i}$ is the mean value, and $\sigma_{i}$ is the standard deviation). Different Gaussian peaks correspond to the charge responses for different numbers of PEs. As shown in Figure 11, by fitting the charge spectrum with a multi-Gaussian function, the gain is determined from the separation between adjacent peaks.
$\begin{split}f=\sum_{i=1}^{n}N_{i}\cdot\text{Gaus}(\mu_{i},\sigma_{i})\end{split}$ (5.1)
Figure 11: Multi-Gaussian Function Fit to Charge Distribution. The horizontal axis is measured in ADC counts, where 1.8 volts correspond to $2^{16}$ ADC counts. Each peak represents the charge response for a specific number of PE: the first peak corresponds to 1PE, the second to 2 PEs, and so on. The separation between adjacent PE peaks indicates the gain.
The gain calibration result is shown in Figure 12. The unit of gain is ADC counts, where 1.8 volts corresponds to $2^{16}$ ADC counts. The bias is 0.084% and the standard deviation is 0.094%.
Figure 12: Gain Calibration Results. The left panel: comparison between the calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with the each bin value from the left 2D histogram.
5.2 Internal Optical Crosstalk
There are two methods for calibrating the IOCT in the TAO CD: multiple PEs hit analysis and generalized Poisson fitting. For multiple PEs hit analysis, the calibrated quantity $P_{\rm IOCT}$ represents the probability that a single electrical pulse generates IOCT pulses. In generalized Poisson fitting, the calibrated quantity $\lambda_{\rm IOCT}$ is the Poisson parameter of the IOCT. According to the mathematical properties of the Generalized Poisson distribution [41], the conversion relationship between these two quantities is given by Equation 5.2.
$P_{\rm IOCT}=1-e^{-\lambda_{\rm IOCT}}$ (5.2)
5.2.1 Multiple Photoelectrons Hit Analysis
Hits occurring before the trigger time are generated by dark noise pulses and EOCT pulses induced by dark noise, which are referred to as equivalent dark noise in this paper. These equivalent dark noise pulses can induce IOCT. When IOCT is induced by these pulses, a multiple PEs hit composed of both the equivalent dark noise and IOCT signals is invariably generated. In contrast, the probability of forming a multiple PEs hit from equivalent dark noise alone is minimal. As shown in Figure 13, it can be observed that multiple PEs hits are primarily contributed by the IOCT effect.
Figure 13: Charge Spectrum Distribution with and without IOCT. The red line represents a case with IOCT ($\lambda_{\rm IOCT}$ = 0.15), and the blue line represents a case without IOCT. It can be observed that the multiple PEs component is mainly contributed by IOCT.
Therefore, the multiple PEs region of the charge spectrum acquired in a dark environment is dominated by IOCT. Hence, the IOCT rate can be quantified using Equation 5.3, where $P_{\rm IOCT}$ is the probability of IOCT occurrence, $N_{\rm multPE}$ is the number of multiple PEs hits and $N_{\rm total}$ is the total number of hits.
$P_{\rm IOCT}=\frac{N_{\rm multPE}}{N_{\rm total}}$ (5.3)
Figure 14: IOCT Calibration Results Based on Multiple Photoelectrons Hit Analysis. The left panel: comparison between the calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with the each bin value from the left 2D histogram.
The IOCT calibration result is shown in Figure 14, the bias is 5.70% and the standard deviation is 2.73%. The bias in the calibration from multiple PEs counting contributions caused by effects other than IOCT. By enabling and disabling various SiPM and electronics effects in the TAO simulation, we found that the multiple PEs hits excluding those from IOCT are dominated by two components: the multiple PEs hits generated by the equivalent dark noise itself, and contributions from the AP. The calibration bias introduced by the multiple PEs contribution of the equivalent dark noise itself is 4.57%, and the bias from the AP is 1.42%. It is clear that the IOCT calibration bias is dominated by the contribution of multiple PEs equivalent dark noise, which necessitates a dedicated correction scheme to mitigate and eliminate this bias.
5.2.2 Generalized Poisson Fitting
To eliminate the influence of multiple PEs equivalent dark noise on the IOCT calibration, the charge spectrum can be fitted using a generalized Poisson distribution. The total distribution of the number of pulses from the $\mu$ Poisson primary pulses and IOCT follows a the generalized Poisson distribution. Since IOCT pulses have no delay relative to the primary pulses, the distribution of the number of pulses corresponds to the number of PEs.
For the equivalent dark noise signal, $\mu=\text{DCR}^{*}\cdot t_{W}\cdot S$, where $\text{DCR}^{*}$ is equivalent DCR, $t_{W}$ is the average width of a single PE waveform, and S is the sensitive area of the SiPM in one channel. By replacing N in Equation 5.1 with GP in Equation 3.3 and fitting this function to the charge spectrum (as shown in Figure 15), the fitted value of $\lambda_{\rm IOCT}$ corresponds to the calibration result of IOCT.
The IOCT calibration result is shown in Figure 16, the bias is 1.40% and the standard deviation is 3.02%. This result indicates that the bias induced by multiple PEs equivalent dark noise has been eliminated.
Figure 15: Generalized Poisson Distribution Fit to the Charge Spectrum.
Figure 16: IOCT Calibration Results Based on Generalized Poisson Fitting. The left panel: comparison between the calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with the each bin value from the left 2D histogram.
6 External Optical Crosstalk Calibration
This paper presents a methodology for calibrating the EOCT rate and the distribution of EOCT emission angle in the TAO CD by switching the SiPM high voltage and using an LED external trigger to acquire data.
As shown in Equation 6.1, when all SiPMs in the TAO CD are operational, the electrical pulses generated within a SiPM include PEs produced by the LED, dark noise, IOCT, AP, and EOCT from other SiPMs. Therefore, the charge $Q_{O}$ consists of contributions from the LED photon signal ($Q_{\rm LED}$), dark noise ($Q_{\rm DN}$), IOCT ($Q_{\rm IOCT}$), AP ($Q_{\rm AP}$) and EOCT ($Q_{\rm EOCT}$). When other SiPMs are turned off, the charge $Q_{C}$ only includes LED, dark noise, IOCT, and AP components. Thus, when acquiring "turn on" and "turn off" data over the same time duration, the difference between $Q_{O}$ and $Q_{C}$ corresponds to the contribution of EOCT. Here, the EOCT ratio is defined as $Q_{\rm EOCT}$ divided by $Q_{C}$.
$\begin{split}Q_{O}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}+Q_{\rm EOCT}\\ Q_{C}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}\\ \text{Rate}_{\rm EOCT}&=\frac{Q_{\rm EOCT}}{Q_{C}}\\ &=\frac{Q_{O}-Q_{C}}{Q_{C}}\end{split}$ (6.1)
In the TAO CD, the operating voltage of each SiPM is controlled via the high-voltage system. Specifically, an SiPM can be turned off by setting its voltage below the breakdown voltage. Since the number of active SiPMs in the CD varies before and after SiPM switching, data can not be acquired using the nhit self-trigger. Therefore, an LED external trigger system is employed. Since high voltage is applied to a single SiPM tile, the EOCT calibration in this study can only provide tile-level calibration results rather than channel-level results. (In TAO CD, each tile corresponds to two channels)
6.1 External Optical Crosstalk Rate Calibration
Three LED external trigger data sets were simulated to study the SiPM EOCT rate calibration. The first data set had all SiPMs turned on; the second had all SiPMs turned on but with the EOCT effect disabled in the simulation; the third had only one SiPM tile turned on. The EOCT rate calibration result can be derived from the first and third data sets, and the true EOCT rate in the simulation can be obtained from the first and second data sets. By comparing the calibration with the true value, the bias and standard deviation can be determined.
The EOCT rate calibration result is shown in Figure 17, the bias is 0.009% and the standard deviation is 0.096%.
Figure 17: EOCT calibration results based on the SiPM on-off switching method. The upper panel: SiPM-level EOCT calibration results. The middle panel: comparison between the calibration results and the simulation truth in each SiPM. The lower panel: 1D histogram filled with each bin value from the middle 2D histogram.
The bias and standard deviation are small because the calibrated EOCT rate (defined in Equation 6.1) is a directly observable quantity rather than a physical parameter such as the EOCT probability for each SiPM. The bias is quantified as the relative difference between the rates from Equation 6.1 and Equation 6.2 (In Equation 6.2, $\text{Rate}^{\rm truth}_{\rm EOCT}$ is the true EOCT ratio, $Q_{\rm EOCT}$ is the charge from EOCT, $Q_{C}$ is the charge measured when other SiPMs are turned off). Due to the LED sources can only be deployed along the central axis, the EOCT emission angle distribution must be calibrated to infer the response off the central axis.
$\text{Rate}^{\rm truth}_{\rm EOCT}=\frac{Q_{\rm EOCT}}{Q_{C}}$ (6.2)
6.2 External Optical Crosstalk Emission Angle Distribution Calibration
In this study, the distribution of the SiPM EOCT emission angle is calibrated by using a reference SiPM to detect and quantify EOCT from other SiPMs positioned at different 5-degree angular intervals relative to the reference SiPM. The procedure is as follows: First, both the reference SiPM and the SiPMs within each 5-degree angular interval are turned on, and the total charge $Q_{O}$ on the reference SiPM is recorded. Next, only the reference SiPM is turned on, and data are acquired over the same duration to measure the total charge $Q_{C}$ on the reference SiPM. Finally, the integrated EOCT response ($E(\theta)$) for each angular interval is obtained by subtracting $Q_{C}$ from $Q_{O}$.
According to the mean value theorem for integral, the EOCT response $\epsilon(\theta)$ in each 5-degree angular interval is given by Equation 6.3, where $\Delta\vec{S}$ is the area vector of each SiPM (with magnitude equals to the sensitive area and direction normal to the surface), $\vec{n}$ is the unit direction vector pointing from the emitting SiPM to the reference SiPM.
$\begin{split}\epsilon(\theta)&=\frac{E(\theta)}{\sum\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}\end{split}$ (6.3)
The reference SiPM selected for this study is located at the detector’s equator. This choice was made because apertures are present at the northern and southern poles of the TAO CD. To maximize the collection of EOCT photons, it is essential to position the reference SiPM away from these apertures. Therefore, the equatorial location was determined to be the optimal position.
The EOCT emission angle distribution calibration result is shown in Figure 18. The left part of Figure 18 presents the result when optical reflection on the SiPM surface is disabled in the simulation. In this case, the bias for most angular bins falls within the statistical uncertainties, except for the first bin. Since the SiPMs do not have a 100% fill factor and the EOCT emission angle distribution exhibits its most significant variation within the first bin, the average value over the sensitive area differs from that over the total area in this bin. The right part displays the result with optical reflection on the SiPM surface enabled. Here, discrepancies between the measured and true values are observed at large angles. However, as the EOCT photon distribution is predominantly concentrated at small angles, the impact of large-angle deviations on the experiment remains acceptable.
Figure 18: The EOCT Emission Angle Distribution Calibration Result. The two figures on the left display result without the SiPM surface reflection effect, while the two figures on the right show result including the reflection effect. The upper row uses a linear y-axis to visualize result in the small-angle range, and the lower row employs a logarithmic y-axis to present result in the large-angle range.
Since the LED light field distribution may be non-uniform, this study simulated a non-uniform LED light field distribution, as illustrated in Figure 19. The corresponding EOCT emission angle distribution calibration result is shown in the left part of Figure 20. This result indicates that the non-uniform LED light field distribution affects the EOCT emission angle distribution calibration.
Figure 19: LED Non-Uniform Light Field Distribution Configuration in Simulation. The distribution follows a Gaussian distribution $\text{Gaus}(\theta,90^{\circ},30^{\circ})$, where the parameters correspond to the polar angle values in the spherical coordinate system.
In the TAO calibration system, the $^{68}{\rm Ge}$ source produces uniform scintillation light in the liquid scintillator, allowing us to correct the non-uniformity of LED light field. The charge response of each SiPM for both the $^{68}{\rm Ge}$ source and the LED source are acquired with all SiPMs turned on. During the EOCT emission angle distribution calibration, the charge measured by the reference SiPM in the "turn on" data is multiplied by a correction factor $\alpha$ defined in Equation 6.4. The summation is performed over all SiPMs within the corresponding angular range, $Q_{{\rm Ge}}$ denotes the SiPM charge response of to the $^{68}{\rm Ge}$ source, and $Q_{\rm LED}$ denotes the SiPM charge response of to the LED light source.
Figure 20: The EOCT Emission Angle Distribution Calibration Result by Using a Non-Uniform Light Field. The two figures on the left show result without correction, while the two figures on the right show the result with correction. The upper row uses a linear y-axis to visualize result in the small-angle range, and the lower row uses a logarithmic y-axis to display result in the large-angle range.
$\alpha=\frac{\sum Q_{{\rm Ge}}}{\sum Q_{\rm LED}}$ (6.4)
The corrected result is shown in the right part of Figure 20, where the influence of the non-uniform LED light field has been successfully corrected.
6.3 Correct Dark Count Rate Base on the External Optical Crosstalk Rate Calibration Result
The DCR calibrated in Section 4.1 includes contributions from EOCT, which is the equivalent dark count rate (EDCR). Therefore, a correction is necessary. The correction factor $\beta_{\rm DN}$:
$\begin{split}\text{DCR}&=\text{EDCR}\cdot\beta_{\rm DN}\\ \beta_{\rm DN}&=\frac{Q_{\rm DN}}{Q_{\rm DN}+Q_{\rm DN,\rm EOCT}}\end{split}$ (6.5)
This correction factor is characterized by the ratio of the dark noise charge response $Q_{\rm DN}$ to the total charge response from dark noise and dark noise-induced EOCT $Q_{\rm DN}+Q_{\rm DN,\rm EOCT}$.
Since the AP of the TAO SiPM is very low (<1%) [42], $\text{Rate}_{\rm EOCT}$ (from Equation 6.1) can be expressed as:
$\begin{split}\text{Rate}_{\rm EOCT}&=\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}}\\ &\approx\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}}\\ &=\frac{\delta_{\rm EOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}{Q_{\rm LED}+Q_{\rm DN}+\delta_{\rm IOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}\end{split}$ (6.6)
where $\delta_{\text{EOCT}}$ denotes the mathematical expectation of EOCT generated by the avalanche signal from a photon hit or a dark noise event, and $\delta_{\text{IOCT}}$ denotes the mathematical expectation of IOCT generated by the avalanche signal from a photon hit or a dark noise event. For IOCT, the number of IOCT induced by an avalanche signal follows the Borel distribution. Based on the mathematical properties of the Borel distribution, $\delta_{\text{IOCT}}$ satisfies:
$\begin{split}\delta_{\rm IOCT}&=\frac{\lambda_{\rm IOCT}}{1-\lambda_{\rm IOCT}}\end{split}$ (6.7)
while $\delta_{\text{EOCT}}$ clearly satisfies:
$\begin{split}\delta_{\rm EOCT}&=\frac{Q_{\rm DN,\rm EOCT}}{Q_{\rm DN}}\end{split}$ (6.8)
According to Equation 6.5, Equation 6.6, Equation 6.7 and Equation 6.8, $\beta_{\rm DN}$ can be expressed as:
$\begin{split}\beta_{\rm DN}&=\frac{1}{1+\frac{\text{Rate}_{\rm EOCT}}{1-\lambda_{\rm IOCT}}}\end{split}$ (6.9)
The DCR result calibration after correction is shown in Figure 21. The bias is -0.40% and the standard deviation is 1.32%.
Figure 21: DCR calibration results corrected based on EOCT calibration results. The left panel: comparison between calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with each bin value from the left 2D histogram.
During the EOCT calibration process, some data is acquired with only a single SiPM turned on. This data can be used to calibrate the DCR without the influence of the EOCT effect. It provides a reference for independent verification and comparative analysis.
7 Discussion
7.1 Impact of Temperature Fluctuation on the TAO SiPM Parameters
As reported in Ref. [9], the dependence of SiPM parameters on temperature and over voltage (OV) can be parameterized as:
$\begin{split}\lambda_{\rm IOCT}&=k_{\rm IOCT}\cdot\text{OV}\,,\qquad\lambda_{\rm EOCT}=k_{\rm EOCT}\cdot\text{OV}\,,\qquad\text{Gain}=k_{\text{Gain}}\cdot\text{OV}\,\\ \text{PDE}&=k_{\text{PDE}}\cdot(\text{OV}-\text{OV}_{\text{ref}})+\text{PDE}(\text{OV}_{\text{ref}})\\ \text{DCR}&=A_{\text{amp}}\cdot\text{OV}\cdot T^{3/2}\cdot e^{-E_{g}/kT}\ \end{split}$ (7.1)
According to the TAO SiPM burn-in test result [8], the relationship between the SiPM OV and temperature follows:
$\begin{split}\text{OV}=V_{\text{ex}}-V_{\text{bd}}=V_{\text{ex}}-(V_{\text{{bd}},{\text{ref}}}+k_{T}\cdot\Delta T)\end{split}$ (7.2)
[…]
Substituting Equation 7.4 into the expression for $\sigma_{T,\text{DCR}}$ in Equation 7.3, the following expression can be derived:
[…]
Parameter Relative Uncertainty
DCR 0.75%
Gain 0.17%
PDE 0.058%
IOCT 0.17%
EOCT Rate 0.17%
7.2 Impact of SiPM Parameter Uncertainties on the Reconstructed Vertex and Energy
In the TAO simulation, we individually add the uncertainties of each SiPM parameter, which originate from temperature fluctuation and the calibration method (the standard deviation derive from calibration). We simulate 1 MeV kinetic energy electrons at the center of the liquid scintillator in the TAO CD, reconstruct the vertex and energy of each electron via the charge center method [39], and derive the corresponding reconstructed vertex uncertainty and energy resolution for both scenarios with and without the inclusion of SiPM parameter uncertainties. The results are summarized in Table 3.
Table 3: Impact of SiPM Parameter Uncertainties on Reconstructed Vertex Uncertainty and Energy Resolution
SiPM Parameter Uncertainty Source Vertex Uncertainty Energy Resolution
Absolute ($\mathrm{m}\mathrm{m}$) Relative ($\%$) Absolute ($\%$) Relative ($\%$)
Baseline No Uncertainty $13.080$ $0.000$ $1.8081$ $0.0000$
DCR Temperature $13.082$ $0.015$ $1.8084$ $0.0166$
PDE $13.082$ $0.015$ $1.8084$ $0.0166$
Gain $13.082$ $0.015$ $1.8104$ $0.1272$
IOCT $13.081$ $0.008$ $1.8083$ $0.0111$
EOCT $13.082$ $0.015$ $1.8086$ $0.0277$
All Parameters $13.089$ $0.0688$ $1.8117$ $0.1991$
DCR Calibration $13.083$ $0.023$ $1.8085$ $0.0221$
PDE $13.099$ $0.145$ $1.8107$ $0.1438$
Gain $13.081$ $0.008$ $1.8089$ $0.0442$
IOCT $13.097$ $0.130$ $1.8134$ $0.2931$
EOCT $13.082$ $0.015$ $1.8084$ $0.0166$
All Parameters $13.121$ $0.313$ $1.8270$ $1.0453$
All Parameters All Uncertainties $13.131$ $0.390$ $1.8300$ $1.2112$
According to the results in Table 3, it can be seen that the SiPM parameter uncertainties induced by temperature fluctuation and those introduced by the calibration method have a very limited impact on the vertex and energy reconstruction precision obtained via the charge center method [39]. The additional vertex smearing and extra degradation of energy resolution introduced by these uncertainties are very small, which directly confirms that the calibration precision fully meets the performance requirements of the TAO detector.
8 Conclusion
This work presents a systematic and comprehensive calibration strategy for the SiPMs in the TAO CD, covering the core performance parameters including DCR, time offset, relative PDE, gain, and IOCT rate. Furthermore, we propose a novel method based on the SiPM on-off switching scheme, which is specifically designed for the calibration of the EOCT rate and its emission angular distribution in detector with high dark noise levels. The feasibility of all proposed calibration methods is fully verified based on the TAO offline software simulation framework, with the bias, standard deviation, and underlying physical sources of bias for each method systematically quantified.
For the key SiPM parameters, the core calibration results are summarized as follows: the DCR calibration achieves a bias of -0.4% and a standard deviation of 1.32% after correction with the EOCT rate calibration results, which effectively suppresses the 23.6% large bias caused by EOCT; the time offset calibration yields both bias and standard deviation better than 0.2 ns, where the uncertainty is dominated by the approximately 1 cm positional deviation of the deployed source inside the detector; for the relative PDE calibration, after correcting for the contribution of dark noise, the method achieves a maximum bias of approximately 3% and a standard deviation of 0.17%, which originates from the non-uniformity of the light field induced by optical reflection on the SiPM surface; the IOCT rate calibration using the Generalized Poisson distribution fitting method achieves a bias of 1.40% and a standard deviation of 3.02%, which resolves the 5.70% bias from dark noise-induced multiple PEs hits, with the residual bias dominated by the contribution from AP.
For the proposed novel EOCT calibration method, the SiPM on-off switching scheme enables EOCT rate calibration with both bias and standard deviation better than 0.1%. Meanwhile, we realize the calibration of the EOCT emission angle distribution by controlling the SiPMs on-off in each angular interval, achieving a bias better than 4% and a standard deviation better than 1% at a 5-degree angular interval.
In addition, this work investigates the SiPM parameter uncertainties induced by temperature fluctuation in the TAO CD. We incorporate the uncertainties from both temperature fluctuation and the calibration method into the TAO simulation, and evaluate their impact on event reconstruction via the charge center method. The results demonstrate that the corresponding impact is very limited, which confirms that the proposed calibration scheme fully meets the stringent performance requirements of the TAO detector.
This work provides solid theoretical guidance for the calibration of individual SiPM performance parameters in the future TAO real detector. Accurate calibration of SiPM parameters will facilitate an in-depth understanding of the TAO detector response and the origins of its energy resolution, and further support the optimization of particle event reconstruction algorithms and the overall performance of the detector.
Acknowledgments
We are gratefully for the JUNO-TAO and the technical staff of the participating institutions. This work is supported by the National Key Research and Development Project of China, Grant No. 2022YFA1602002, National Natural Science Foundation of China under Grant Number 12275281.